### v1

# Early stopping for conjugate gradients in statistical inverse problems

Abstract
We consider estimators obtained by iterates of the conjugate gradient (CG) algorithm applied to the normal equation of prototypical statistical inverse problems. Stopping the CG algorithm early induces regularisation, and optimal convergence rates of prediction and reconstruction error are established in wide generality for an ideal oracle stopping time. Based on this insight, a fully data-driven early stopping rule $\tau$ is constructed, which also attains optimal rates, provided the error in estimating the noise level is not dominant.
The error analysis of CG under statistical noise is subtle due to its nonlinear dependence on the observations. We provide an explicit error decomposition into two terms, which shares important properties of the classical bias-variance decomposition. Together with a continuous interpolation between CG iterates, this paves the way for a comprehensive error analysis of early stopping. In particular, a general oracle-type inequality is proved for the prediction error at $\tau$. For bounding the reconstruction error, a more refined probabilistic analysis, based on concentration of self-normalised Gaussian processes, is developed. The methodology also provides some new insights into early stopping for CG in deterministic inverse problems. A numerical study for standard examples shows good results in practice for early stopping at $\tau$.
Keywords: linear inverse problems, conjugate gradients, partial least squares, early stopping, discrepancy principle, adaptive estimation, oracle inequalities
MSC Classification: 65J20, 65F10, 62G05
1 Introduction
The conjugate gradient (CG) algorithm is arguably one of the computationally most efficient off-the-shelf methods for solving systems of linear equations. Like for standard gradient descent methods, stopping the algorithm early, that is before terminating at the solution, induces a regularisation. In the context of deterministic ill-posed inverse problems the seminal work by Nemirovskii [20] and Hanke [10] has shown that a stopping criterion based on the discrepancy principle can lead to optimal convergence rates under bounded noise perturbations. Despite its importance in applications, the case of statistical noise, that is random perturbations, has been understood only partially for CG so far. Major reasons are the nonlinearity of the CG algorithm and that canonical Gaussian white noise $\xi$ on an infinite-dimensional Hilbert space has infinite norm $\lVert\xi\rVert=\infty$ almost surely. Therefore, the discrepancy principle cannot be well-defined.
One way out is to modify the stopping criterion so that the noise in the residual is smoothed out. Blanchard and Mathé [6] build on this idea and use Gaussian concentration together with the deterministic CG analysis to derive convergence rates, which are, however, suboptimal in terms of a-posteriori error control. In the more involved kernel learning setting a similar approach is taken by Blanchard and Krämer [5], defining the CG algorithm in the underlying energy norm, which regularises the noise. The main point there is to establish good convergence properties of the best iterate along the entire iteration path and then leave the choice of the iterate to a-posteriori model selection on all iterates. An interesting extension to time series data is provided by Singer, Krivobokova and Munk [22]. In the statistics and machine learning community CG-type methods are very popular under the name partial least squares, see the excellent survey by Rosipal and Krämer [21] and the recent paper by Finocchio and Krivobokova [9].
Our approach to conjugate gradients for statistical inverse problems is intrinsically different. We consider the standard CGNE (conjugate gradients for the normal equation) algorithm, which when implemented is necessarily finite-dimensional, and analyse it carefully, keeping explicitly track of the underlying dimension $D$. This is in line with the approach by Blanchard, Hoffmann and Reiß [3, 4] for linear spectral methods, which has been compared to smoothed residual approaches by Stankewitz [23]. Due to the nonlinear dependence on the noise, however, the analysis must be undertaken for every noise realisation and often requires more sophisticated arguments. This precise nonasymptotic analysis provides also for unspecified perturbations $\xi$ in the data new structural insights for the CGNE algorithm applied to inverse problems. Throughout the text a deterministic noise and a Gaussian white noise model are considered as concrete examples. The deterministic noise results are essentially known in the literature but obtained differently and thus serve as an easy proof of concept. The main contributions of this work, when specified to the Gaussian model, are the following:
  • Introducing interpolation between CG iterates, we prevent overshooting and avoid the analysis of update polynomials.
  • We identify two random quantities, called stochastic error and approximation error, which share important properties of variance and bias and allow for a precise nonasymptotic CG error control.
  • Equilibrating stochastic and approximation error, an oracle stopping time achieves optimal prediction error control under minimal assumptions on the unknown signal and on the noise. In particular, under mild singular value conditions, the best CG iterate will be as good as the best iterate in the prototypical Showalter method, that is standard gradient descent flow.
  • We construct a data-driven stopping rule that satisfies an oracle-type inequality for the prediction error and achieves optimal convergence rates whenever the error in estimating the noise level $\lVert\xi\rVert^{2}$ is not dominant.
  • The convergence rates transfer to the reconstruction error, thus establishing minimax optimality of the best iterate along the CG iteration path and – under restrictions on the dimension $D$ – of the iterate selected by our early stopping criterion.
In the following Section 2, the setting is introduced. Section 3 explains the main ideas and results of the paper. Afterwards, we progress in detail. Results for the interpolated CG iterates are presented in Section 4, and the prediction error is analysed in Section 5. Section 6 derives our stopping criterion and proves an oracle-type inequality for its prediction error. The results are transferred in Section 7 to the reconstruction error. The numerical results for some standard examples in Section 8 are in good agreement with the theory. Technical results and proofs are delegated to Section 9.
2 The setting
We consider the classical prototypical model for statistical inverse problems, see e.g. Cavalier [7] for a survey. We observe
$Y=g+\xi=Af+\xi,$ (2.1)
where $A:H_{1}\to H_{2}$ is a known bounded linear operator between real Hilbert spaces $H_{1}$ and $H_{2}$, $f\in H_{1}$ is the unknown signal of interest and $\xi$ is an error term. In practice, the problem is always discretised such that we assume $H_{1}={\mathbb{R}}^{P}$, $H_{2}={\mathbb{R}}^{D}$ with $D\geqslant 3$ and $\xi$ to be ${\mathbb{R}}^{D}$-valued. Suppose that $A:{\mathbb{R}}^{P}\to{\mathbb{R}}^{D}$ is surjective and thus $\operatorname{rank}(A)=D\leqslant P$, which can always be achieved by restricting the observations $Y$ to the range of $A$. We wish to recover the minimum-norm least-squares solution $f^{\dagger}\coloneqq A^{\dagger}g=A^{\top}(AA^{\top})^{-1}g$ of $g=Af$ from the observation $Y$ by applying the CG iterates as defined in Section 3.
For the analysis (not the algorithm), the model (2.1) can be transformed using the singular value decomposition (SVD) of $A$: Let $\lambda_{1}\geqslant\dots\geqslant\lambda_{D}>0$ be the nonzero singular values of $A$. Denote by $u_{1},\dots,u_{D}$ and $v_{1},\dots,v_{P}$ the corresponding left- and right-singular vectors, respectively. Then we obtain the equivalent sequence space model [7, Eq. (1.5)]
$Y_{i}=g_{i}+\xi_{i}=\lambda_{i}f_{i}+\xi_{i},\quad i=1,\dots,D,$
[…]
When specifying convergence rates, we will consider signals $f$ satisfying the Sobolev-type source condition
$\lVert f\rVert_{\mu}\coloneqq\Big{(}\sum_{i=1}^{D}\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/2}\leqslant R$
for $\mu>0$, $R\geqslant 1$. Note that the assumption $\lVert f\rVert_{\mu}\leqslant R$ on $f$ corresponds to $\lVert g\rVert_{\mu+1/2}\leqslant R$ on $g=Af$.
For some results we will suppose that the inverse problem is mildly ill-posed with singular values $(\lambda_{i})$ satisfying the polynomial spectral decay assumption
$c_{A}i^{-p}\leqslant\lambda_{i}\leqslant C_{A}i^{-p},\quad i=1,\dots,D,$ (PSD($p$,$c_{A}$,$C_{A}$))
where $p\geqslant 0$ is fixed and $0<c_{A}\leqslant C_{A}<\infty$ are numerical constants.
Throughout the paper, (in-)equalities up to absolute constants are denoted by $\cong$’, $\lesssim$’, $\gtrsim$’. Further dependencies on the constants will be highlighted by additional indices. Bounds of the type $B_{1}\lesssim_{c_{A}}B_{2}$, $B_{3}\lesssim_{C_{A}}B_{4}$ indicate that they hold true already under a one-sided condition, i.e. for $C_{A}=\infty$ and $c_{A}=0$, respectively.
First results will be obtained for any error $\xi$ or under minor assumptions, see e.g. the minimax bounds of Theorem 5.8 in conjunction with Remark 5.10. To analyse early stopping under random noise later, we often specify to the Gaussian noise model
$\xi\coloneqq\delta Z\quad\text{with}\quad Z\sim\mathcal{N}(0,I_{D})\quad\text{and}\quad\delta\in(0,1),$ (GN)
where we denote by $I_{D}$ the identity matrix in ${\mathbb{R}}^{D\times D}$. In contrast, in the classical setting of deterministic inverse problems, we study noise $\xi$ with
$\lVert Y-Af\rVert=\lVert\xi\rVert\leqslant\delta\quad\text{and}\quad\delta\in(0,1).$ (DN)
Alternatively, if $\xi\sim\mathcal{N}(0,\Sigma)$ with $\operatorname{tr}(\Sigma)\leqslant\delta^{2}$, then $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]\leqslant\delta^{2}$ and the results for (DN) will hold in a similar manner by taking expectations.
To avoid special cases, we will suppose throughout the paper that
$\sum_{j=1,\dots,D:\,\lambda_{j}=\lambda_{i}}Y_{j}^{2}>0,\quad i=1,\dots,D.$ (Y)
Under (GN), this is almost surely the case. See Remark 5.4 for a discussion of this assumption.
Let us introduce some further notation. Define
$\lVert\mathsf{p}\rVert_{Y}\coloneqq\lVert p(AA^{\top})Y\rVert,\quad\langle{\mathsf{p}},{\mathsf{q}}\rangle_{Y}\coloneqq\langle{p(AA^{\top})Y},{q(AA^{\top})Y}\rangle$
for polynomials $p,q$ in the set $\operatorname{Pol}_{k}$ of real polynomials with maximal degree $k$. By Assumption (Y), $\lVert\mathsf{p}\rVert_{Y}$ is indeed a norm and $\langle{\mathsf{p}},{\mathsf{q}}\rangle_{Y}$ an inner product on $\operatorname{Pol}_{k}$ for $k<d$. For $k=d$ both lack positive definiteness. Consider the linear and affine subspaces
$\operatorname{Pol}_{k,0}\coloneqq\big{\{}p\in\operatorname{Pol}_{k}\,\big{|}\,p(0)=0\big{\}},\quad\operatorname{Pol}_{k,1}\coloneqq\big{\{}p\in\operatorname{Pol}_{k}\,\big{|}\,p(0)=1\big{\}}.$
We often abbreviate $h(AA^{\top})v$ for functions $h$ and $v\in{\mathbb{R}}^{D}$ by $\mathsf{h}v$. By abuse of notation, we sometimes write the value of a function $h$ evaluated at $x$ instead of the function itself. For example, $\lVert\mathsf{x^{1/2}}v\rVert=\lVert\mathsf{h}v\rVert=\lVert(AA^{\top})^{1/2}v\rVert$ with $h(x)=x^{1/2}$, $x\geqslant 0$.
Denote by $\delta_{l}$ the Dirac measure in $l\in{\mathbb{R}}$ and put $\delta_{lk}\coloneqq\delta_{l}(\{k\})$ for $l,k\in{\mathbb{R}}$. We set $a\wedge b=\min(a,b)$, $a\vee b=\max(a,b)$ and $a_{+}=a\vee 0$ for $a,b\in{\mathbb{R}}$.
3 Overview of main results
The original CG method [13] is used for numerically solving systems of linear equations, where the coefficient matrix is symmetric and positive (semi-)definite. This is not necessarily the case in our model. Following the standard technique [10, Section 2.3], we apply the iterative CG algorithm to the normal equation $A^{\top}A\widehat{f}=A^{\top}Y$ [2, Algorithm 7.4.1] and choose the initial value $\widehat{f}_{0}=0$. The iterates $\widehat{f}_{k}$, for $k=0,\dots,d$ or up to some stopping criterion, yield our estimators of $f$. For the theoretical analysis in this paper, the estimators are defined via minimising residual polynomials, which emerge from the theory of Krylov subspace methods [10, Section 2].
Definition 3.1.
At iteration $k=0,\dots,d$, the prediction estimator $\widehat{g}_{k}$ of $g=Af$ and the CG estimator $\widehat{f}_{k}$ of the minimum-norm solution $f^{\dagger}$ of $g=Af$ are given by
$\widehat{g}_{k}\coloneqq(I_{D}-r_{k}(AA^{\top}))Y\quad\text{and}\quad\widehat{f}_{k}\coloneqq A^{\dagger}\widehat{g}_{k},$
where the residual polynomial $r_{k}$ is obtained as
$r_{k}\coloneqq\operatorname*{argmin}_{p_{k}\in\operatorname{Pol}_{k,1}}\lVert p_{k}(AA^{\top})Y\rVert^{2}.$
In the case $k=d$, $r_{d}$ maps all squared nonzero singular values of $A$ to zero implying $\widehat{f}_{d}=A^{\dagger}Y$. Thus, the CG estimator at iteration $d$ is a minimum-norm solution of (2.1) and $d$ is the maximal iteration index.
One new idea is to allow also for noninteger iteration indices $t\in[0,d]$ by interpolating linearly between the residual polynomials and thus also between the estimators. The time-continuous residual polynomials have useful properties, which are discussed in Section 4 in detail.
Since the estimator $\widehat{g}_{t}$ depends nonlinearly on the observation $Y$, we cannot calculate a standard bias-variance decomposition, but need to analyse the prediction error for any realisation of $\xi$. In Proposition 5.2, we are able to decompose the error nonasymptotically into terms that share important properties of stochastic and approximation errors, see Lemma 5.3 for a detailed analysis.
Corollary 3.2.
For $t\in[0,d]$, the prediction error satisfies
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2(A_{t,\lambda}+S_{t,\lambda}),$
where we have for the stochastic error term $S_{t,\lambda}$ that
  1. (a)
    $S_{t,\lambda}=0$ for $\xi=0$,
  2. (b)
    $S_{0,\lambda}=0$ and $S_{d,\lambda}=\lVert\xi\rVert^{2}$,
  3. (c)
    $S_{t,\lambda}$ is nondecreasing in $t$,
and for the approximation error term $A_{t,\lambda}$ that
  1. (a)
    $A_{t,\lambda}\leqslant 0$ if $Af=0$,
  2. (b)
    $A_{0,\lambda}=\lVert Af\rVert^{2}$ and $A_{d,\lambda}=0$,
  3. (c)
    $A_{t,\lambda}$ has a nonnegative upper bound given in 5.3(h), which is nonincreasing in $t$.
Both error terms, defined in Definition 5.1, depend on the observation through the residual polynomial. In addition, the stochastic error depends on the noise $\xi$ and the approximation error on the signal $g=Af$. Therefore, they are unknown to the user, but they play a key role in the nonasymptotic analysis of early stopping. Given the monotonicity properties of $S_{t,\lambda}$ and the upper bound of $A_{t,\lambda}$, it is natural to consider the following sequential oracle index.
Definition 3.3.
The (data-dependent) balanced oracle index is given by
$\tau_{\mathfrak{b}}\coloneqq\inf\{t\in[0,d]\,|\,A_{t,\lambda}\leqslant S_{t,\lambda}\}.$
Due to the continuity of the interpolations $A_{t,\lambda}$ and $S_{t,\lambda}$ in $t$ and $A_{0,\lambda}\geqslant S_{0,\lambda}$, $A_{d,\lambda}\leqslant S_{d,\lambda}$, the approximation and stochastic error are indeed balanced at $\tau_{\mathfrak{b}}$, meaning that $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$.
The choice is motivated by the fact that the bound $2(A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda})$ of the prediction error at the balanced oracle is by the monotonicity of $t\mapsto S_{t,\lambda}$ at most twice as large as the minimal prediction error bound:
$A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}=2S_{\tau_{\mathfrak{b}},\lambda}\leqslant 2\inf_{t\in[0,d]}(A_{t,\lambda}\vee S_{t,\lambda})\leqslant\inf_{t\in[0,d]}2(A_{t,\lambda}+S_{t,\lambda}).$
The balanced oracle $\tau_{\mathfrak{b}}$ serves as a theoretical benchmark, not accessible to the user, that we shall try to mimic in a data-driven way, see Tsybakov [26, Section 1.8] for a discussion of the oracle concept.
Assume (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$ and Gaussian noise (GN). In Theorem 5.8, we show that the expected stochastic error $\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]$ and the prediction estimator $\widehat{g}_{\tau_{\mathfrak{b}}}$ at the balanced oracle attain the risk (mean squared error) rate
$\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)\coloneqq R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}$
over Sobolev-type ellipsoids, i.e.
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}=\sup_{g:\,\lVert g\rVert_{\mu+1/2}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{g}_{\tau_{\mathfrak{b}}}-g\rVert^{2}\big{]}\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta).$ (3.1)
From an asymptotic point of view, where the dimension $D=D(\delta)\to\infty$ as $\delta\to 0$, this rate is minimax optimal if $D(\delta)$ grows faster than $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$ as shown in Proposition 5.11. The proof of the upper bound is very general and allows extensions to other source conditions and noise specifications (see Remark 5.10). Also it enables us to show close parallels to Showalter’s method, i.e. standard gradient flow (see Remark 5.13).
For deterministic noise of level $\delta$, the prediction error at the least-squares solution (i.e. at the last iteration $\widehat{f}_{d}$) is $\delta^{2}$ and no regularisation is required. Under (GN), however, this becomes $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{d}-f)\rVert^{2}]=D\delta^{2}$ and regularisation by early stopping is essential for larger dimensions $D$ exactly as in nonparametric regression. It is the first time that the CGNE iteration path is shown to achieve the optimal prediction error rate under (GN) at some iterate.
To define a data-driven stopping rule, we monitor the squared residual norm (also called discrepancy, data misfit or training error)
$R_{t}^{2}\coloneqq\lVert A\widehat{f}_{t}-Y\rVert^{2}=\lVert r_{t}(AA^{\top})Y\rVert^{2}$
at each iteration step, which is greedily minimised in the CG construction. Similarly to Blanchard, Hoffmann and Reiß [3, 4], but for a random balanced oracle $\tau_{\mathfrak{b}}$, we mimic the balanced oracle by a discrepancy-type principle.
Definition 3.4.
For a suitable critical value $\kappa>0$, introduce the early stopping rule
$\tau\coloneqq\inf\big{\{}t\in[0,d]\,\big{|}\,R_{t}^{2}\leqslant\kappa\big{\}}.$
Under (DN), setting $\kappa\coloneqq c^{2}\delta^{2}$ for an a-priori fixed constant $c>1$, this coincides with the well-known discrepancy principle [8, Sections 4.3 and 7.3]. The balanced oracle can also be expressed in terms of the residual norm, and in Remark 6.2, the relation between the stopping indices $\tau$ and $\tau_{\mathfrak{b}}$ is discussed.
Theorem 6.9 yields for any signal $f$ an oracle-type inequality under Gaussian noise (GN), relating the prediction error at $\tau$ to the expected minimal prediction error bound:
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\big{]}\lesssim\inf_{t\in[0,d]}\operatorname{{\mathbb{E}}}[A_{t,\lambda}+S_{t,\lambda}]+\delta^{2}\sqrt{D}+\lvert\kappa-\delta^{2}D\rvert.$ (3.2)
Note that we do not have a classical oracle inequality in terms of the oracle error $\inf_{t\in[0,d]}\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{t}-f)\rVert^{2}]$, since the decomposition into the two error terms in Corollary 3.2 gives only an upper bound. The dimension-dependent error term $\delta^{2}\sqrt{D}$ is due to the variability of the Gaussian noise $\xi$ and unavoidable in the linear case of truncated SVD estimation [3]. Therefore, the early stopped estimator $\widehat{g}_{\tau}$ has the same error rate as $\widehat{g}_{\tau_{\mathfrak{b}}}$ if the remainder term is small enough. In this case of a-posteriori error control, we say that we can adapt to the balanced oracle. In practice, the noise level $\delta$ needs to be estimated, and the threshold $\kappa$ is allowed to deviate from $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D$.
The nonasymptotic oracle inequality (3.2) does not require to fix any source conditions on $f$, yet allows a straight-forward derivation of adaptive asymptotic error bounds, see Cavalier [7, Section 1.3] for more details on this approach. Assume again (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$ and (GN). If the remainder term $\delta^{2}\sqrt{D}+\lvert\kappa-\delta^{2}D\rvert$ is at most of the order of the optimal rate, the bound (3.1) directly carries over such that
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\big{]}=\sup_{g:\,\lVert g\rVert_{\mu+1/2}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{g}_{\tau}-g\rVert^{2}\big{]}\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta),$
and the optimal rate of convergence is attained. The early stopping rule $\tau$ is minimax adaptive for the prediction error over Sobolev-type ellipsoids for all regularities $\mu>0$ with $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$ (see Corollary 6.10).
In Section 7, we transfer the results for the prediction error to the reconstruction error, i.e. the risk of the CG estimator. Similarly to the prediction error, the analysis of the reconstruction error relies on an upper bound given in Proposition 7.1, which depends on the error terms and the residual polynomial.
Under (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$ and (GN), combining Proposition 7.1 with an interpolation inequality and $\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ yields
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)\coloneqq R^{(4p+2)/(4\mu p+2p+1)}\delta^{8\mu p/(4\mu p+2p+1)}.$
Under the same growth condition $D\gtrsim_{\mu,p,c_{A}}(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$ as before, this rate is minimax optimal such that $\widehat{f}_{\tau_{\mathfrak{b}}}$ simultaneously attains the optimal rate of convergence for the prediction as well as for the reconstruction error (see Theorem 7.6).
For the reconstruction error at $\tau$, we need to perform a more intricate high probability analysis of the intrinsic regularisation parameter $\lvert r^{\prime}_{\tau}(0)\rvert$ and the error terms at $\tau$ occurring in the upper bound on the reconstruction error. In Theorem 7.10, under the same assumptions as for the prediction error, we show that the early stopping rule $\tau$ is also minimax adaptive for the reconstruction error over signals $f$ satisfying $\lVert f\rVert_{\mu}\leqslant R$ for all regularities $\mu>0$ with $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$. As a byproduct, our analysis yields also the classical convergence rates for CG under deterministic noise stopped by the discrepancy principle, see Proposition 7.12, where some arguments may be more transparent than previous ones.
In Section 8, we consider some numerical examples from the literature. For mildly ill-posed problems, the practical performance of the data-driven stopping rule $\tau$ is quite remarkable, usually very close to the best iterate along the CG path. Even for a severely ill-posed test problem, which is not covered by our theory, the early stopping criterion provides reasonable results.
4 The interpolated conjugate gradient method
We avoid discretisation errors by interpolating in Definition 4.1 linearly between the residual polynomials from Definition 3.1. From a computational perspective, the interpolation does not impose additional computational cost. In practice, we calculate the estimators at integer iteration indices, and as soon as $R_{m}^{2}\leqslant\kappa$ for some $m\in\{0,\dots,d\}$ with $\kappa$ from Definition 3.4, we stop the computation. The realised value of $\tau\in(m-1,m]$ can then be calculated from the last two residuals. Since CG iterates usually converge very quickly, stopping at an interpolated iteration helps to prevent overshooting. For a discussion how the interpolation allows to retrieve an order-optimality result from the theory of deterministic inverse problems see Remark 7.13.
We analyse in Section 9.2 the classical noninterpolated CG method, which forms the basis for the following results. The proofs for this section are given in Section 9.3.
Definition 4.1.
For $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in(0,1]$, the interpolated residual polynomial is defined as
$r_{t}\coloneqq(1-\alpha)r_{k}+\alpha r_{k+1}.$
The interpolated prediction and CG estimators are given by
$\widehat{g}_{t}\coloneqq(I_{D}-r_{t}(AA^{\top}))Y\quad\text{and}\quad\widehat{f}_{t}\coloneqq A^{\dagger}\widehat{g}_{t},$
respectively, i.e.
$\widehat{g}_{t}=(1-\alpha)\widehat{g}_{k}+\alpha\widehat{g}_{k+1}\quad\text{and}\quad\widehat{f}_{t}=(1-\alpha)\widehat{f}_{k}+\alpha\widehat{f}_{k+1}.$
Remark 4.2.
The interpolation follows also naturally by defining $r_{t}$ via penalised minimisation. Let
$r_{k+1,\lambda}\coloneqq\operatorname*{argmin}_{p_{k+1}\in\operatorname{Pol}_{k+1,1}}\Big{(}\lVert\mathsf{p_{k+1}}\rVert_{Y}^{2}+\lambda\alpha_{k+1}(p_{k+1})^{2}\Big{)},$
where $\alpha_{k+1}(p)\in{\mathbb{R}}$ denotes the coefficient in front of $x^{k+1}$ for a polynomial $p(x)$ and $\lambda>0$ is a penalty parameter. Note that $p_{k+1}-r_{k}-\alpha(r_{k+1}-r_{k})$ lies in $\operatorname{Pol}_{k,0}$ for $\alpha=\alpha_{k+1}(p_{k+1})/\alpha_{k+1}(r_{k+1})$ and $r_{k},r_{k+1}\perp_{Y}\operatorname{Pol}_{k,0}$ by Lemma 9.3. By rewriting $p_{k+1}=r_{k}+\alpha(r_{k+1}-r_{k})+q_{k}$ with $q_{k}\in\operatorname{Pol}_{k,0}$ and applying the Pythagorean theorem, the minimisation yields a polynomial $r_{k+1,\lambda}=r_{k}+\alpha_{\lambda}(r_{k+1}-r_{k})$ with
$\alpha_{\lambda}\coloneqq{}$ $\operatorname*{argmin}_{\alpha\in{\mathbb{R}}}\Big{(}\lVert\mathsf{r_{k}}+\alpha(\mathsf{r_{k+1}-r_{k}})\rVert_{Y}^{2}+\lambda\alpha_{k+1}(r_{k+1})^{2}\alpha^{2}\Big{)}$
[…]
where we used that $\langle{\mathsf{r_{k+1}-r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}=0$ by Lemma 9.3. The numerator is just $R_{k}^{2}-R_{k+1}^{2}$, which is strictly positive for $k+1\leqslant d$ so that $\alpha_{\lambda}\in(0,1]$ holds with $\alpha_{\lambda}\downarrow 0$ for $\lambda\uparrow\infty$ and $\alpha_{\lambda}=1$ for $\lambda=0$. This way we recover $r_{t}$ for $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in(0,1]$, as $r_{k+1,\lambda}$ for a suitable penalisation parameter $\lambda>0$.
We have extended Definition 3.1 for $t=1,\dots,d$, keeping the starting values $\widehat{g}_{0}=\widehat{f}_{0}=0$ and the relation $A\widehat{f}_{t}=\widehat{g}_{t}$ for $t\in[0,d]$. Since there are $d$ distinct nonzero singular values of $A$, the $k$-th residual polynomial has degree $k$ for $k=0,\dots,d$. Therefore, $r_{t}$ for $t\in(k,k+1]$ has degree $k+1$ and is an element of $\operatorname{Pol}_{k+1,1}$. The squared residual norms interpolate nonlinearly:
$R_{t}^{2}=\lVert r_{t}(AA^{\top})Y\rVert^{2}=(1-\alpha)^{2}R_{k}^{2}+(1-(1-\alpha)^{2})R_{k+1}^{2}.$
Lemma 4.3.
For $t,s\in[0,d]$, the squared residual norm satisfies
$R_{t\vee s}^{2}\leqslant\langle{\mathsf{r_{t}}},{\mathsf{r_{s}}}\rangle_{Y}\quad\text{and}\quad\lVert A(\widehat{f}_{t}-\widehat{f}_{s})\rVert^{2}\leqslant\lvert R_{t}^{2}-R_{s}^{2}\rvert.$
Both statements hold with equality for integer $t$ and $s$.
Definition 4.4.
Denote by $x_{1,t}$ the smallest zero on $[0,\infty)$ of the $t$-th residual polynomial $r_{t}$, $t\in(0,d]$. For $x\geqslant 0$, introduce the functions
$r_{t,<}(x)\coloneqq r_{t}(x){\bf 1}(x<x_{1,t}),\quad r_{t,>}(x)\coloneqq r_{t}(x){\bf 1}(x>x_{1,t})$
and set $r_{0,<}\coloneqq r_{0}\equiv 1$, $r_{0,>}\equiv 0$. If $x_{1,t}$ is clear from the context, we also write $f_{<}(x)\coloneqq f(x){\bf 1}(x<x_{1,t})$, $\lVert\mathsf{f}\rVert_{Y,<}\coloneqq\lVert\mathsf{f}_{<}\rVert_{Y}$ and $\langle{\mathsf{f}},{\mathsf{g}}\rangle_{Y,<}\coloneqq\langle{\mathsf{f_{<}}},{\mathsf{g_{<}}}\rangle_{Y}$ for any functions $f$ and $g$, analogously for ‘$>$’.
For integer iteration indices, the residual polynomials possess several well-known properties, which are collected in Lemma 9.5. These carry over to the interpolated analogues. We state in the following lemma the most important properties.
Lemma 4.5 (Properties of the residual polynomial).
Let $t=k+\alpha$ with $k=0,\dots,d-1$ and $\alpha\in(0,1]$.
  1. (a)
    For $k=1,\dots,d-1$, $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$, where $\lVert\cdot\rVert_{\mathrm{spec}}$ denotes the spectral norm on ${\mathbb{R}}^{D\times D}$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<\tilde{\lambda}_{d}^{2}<\dots<\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
  2. (b)
    For $\alpha\in(0,1)$, $r_{t}$ has $k+1$ real simple zeros $x_{i,t}$ satisfying $0<x_{i,k+1}<x_{i,t}<x_{i,k}$, $i=1,\dots,k+1$, setting $x_{1,0},x_{k+1,k}\coloneqq\infty$.
  3. (c)
    $r_{t}$ can be written as
    $r_{t}(x)=\prod_{i=1}^{\lceil t\rceil}\Big{(}1-\frac{x}{x_{i,t}}\Big{)},\ x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\quad\text{with}\quad\lvert r_{t}^{\prime}(0)\rvert=\sum_{i=1}^{\lceil t\rceil}x_{i,t}^{-1}.$
  4. (d)
    $r_{t}$ is nonnegative, decreasing and convex on $[0,x_{1,t}]$ and log-concave on $[0,x_{1,t})$.
  5. (e)
    For $t_{j}=k+\alpha_{j}$, $j=1,2$, with $k=0,\dots,d-1$ and $10<\alpha_{1}<\alpha_{2}<$ 1, we have
    $x_{1,t_{2}}<x_{1,t_{1}}<x_{2,t_{2}}<\dots<x_{k,t_{1}}<x_{k+1,t_{2}}<x_{k+1,t_{1}}.$
  6. (f)
    For fixed $x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]$, $t\mapsto r_{t,<}(x)$ is nonincreasing on $[0,d]$.
  7. (g)
    For $k=1,\dots,d$, $i=1,\dots,k$, we have
    $x_{k+1-i,k}\leqslant\lambda_{i}^{2}.$
The following key proposition without interpolation is originally due to Nemirovskii [20, (3.14)], see Lemma 9.6 for details. We extend it to real iteration indices. It allows the restriction to squared singular values in the interval $[0,x_{1,t}]$ for the error analysis.
Proposition 4.6.
The residual polynomial satisfies $\lVert\mathsf{r_{t}}\rVert_{Y}^{2}\leqslant\lVert\mathsf{r_{t,<}^{1/2}}\rVert_{Y}^{2}$ for all $t\in[0,d]$.
On the interval $[0,x_{1,t}]$, the residual polynomial can be controlled in terms of the value of its first derivative $r_{t}^{\prime}(0)$ in zero. This quantity plays an essential role in the CG analysis, as discussed in Remark 5.9.
The first inequality in the lemma below is well-known, dating back to Louis [17, Lemma 4.3.15], and follows from the convexity of the residual polynomial $r_{t}$ on $[0,x_{1,t}]$. The second one is new and results from the log-concavity of $r_{t}$ on the same interval.
Lemma 4.7.
For all $t\in[0,d]$, we have
$(1-\lvert r_{t}^{\prime}(0)\rvert x)_{+}\leqslant r_{t}(x)\leqslant e^{-\lvert r_{t}^{\prime}(0)\rvert x},\quad x\in[0,x_{1,t}].$
Using the upper bound, we get directly the following control for the approximation error $A_{t,\lambda}$ defined below.
Corollary 4.8.
Under the source condition $\lVert g\rVert_{\mu+1/2}\leqslant R$, for all $t\in(0,d]$, we have
$\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}\leqslant R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}.$
5 Analysis of the prediction error
5.1 Decomposition of the prediction error
Due to the nonlinearity of the CG estimator, we cannot completely disentangle data and approximation errors [8, Section 7.3]. We are, however, able to identify terms that share important properties of these errors, in particular regarding monotonicity along the iteration path.
Definition 5.1.
For $t\in[0,d]$, the stochastic error term is defined as
$S_{t,\lambda}\coloneqq\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2}$
and the approximation error term as
$A_{t,\lambda}\coloneqq\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+R_{t}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}.$
Proposition 5.2.
For $t\in[0,d]$, we have
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}=A_{t,\lambda}+S_{t,\lambda}-2\langle{\xi},{\mathsf{r_{t,>}}Y}\rangle\leqslant 2(A_{t,\lambda}+S_{t,\lambda}).$
Proof.
We deduce by direct calculation
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}$ $=\lVert(\mathsf{1-r_{t}})Y-g\rVert^{2}$
$=\lVert\mathsf{r_{t}}Y\rVert^{2}+\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{t}}Y}\rangle$
$=R_{t}^{2}+\lVert\xi\rVert^{2}-2\big{(}\langle{\xi},{\mathsf{r_{t,>}}Y}\rangle+\langle{\xi},{\mathsf{r_{t,<}}Y}\rangle\big{)}$
$=R_{t}^{2}+\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{t,>}}Y}\rangle+\lVert\mathsf{r_{t,<}^{1/2}}(Y-\xi)\rVert^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}-\lVert\mathsf{r_{t,<}^{1/2}}\xi\rVert^{2}$
$=\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+R_{t}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{t,>}}Y}\rangle$
$=A_{t,\lambda}+S_{t,\lambda}-2\langle{\xi},{\mathsf{r_{t,>}}Y}\rangle.$
When restricting to singular values $\lambda_{i}^{2}\leqslant x_{1,t}$ the same calculation gives
$\lVert A(\widehat{f}_{t}-f)\rVert_{<}^{2}=\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+\lVert\mathsf{r_{t}}Y\rVert_{<}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}.$
On the other hand, we have $\lVert A(\widehat{f}_{t}-f)\rVert_{>}^{2}=\lVert(\mathsf{1-r_{t}})Y-g\rVert_{>}^{2}\leqslant 2(\lVert\xi\rVert_{>}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2})$ and thus
$\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert^{2}$ $\leqslant\lVert A(\widehat{f}_{t}-f)\rVert_{<}^{2}+\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert_{>}^{2}$
$\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+\lVert\mathsf{r_{t}}Y\rVert_{<}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}+\lVert\xi\rVert_{>}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2}$
$=\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+R_{t}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2},$
which equals $A_{t,\lambda}+S_{t,\lambda}$ and gives the upper bound. ∎
Lemma 5.3.
[…]
  1. (d)
    $S_{t,\lambda}\leqslant\lVert((\mathsf{\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$.
[…]
  1. (g)
    $A_{t,\lambda}\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}\leqslant\lVert\mathsf{\exp(-\lvert r_{t}^{\prime}(0)\rvert x/2)}g\rVert^{2}$,
  2. (h)
    the upper bound $\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}$ is nonincreasing in $t$.
Proof.
[…]
Remark 5.4.
For the proof of $S_{d,\lambda}=\lVert\xi\rVert^{2}$ and $A_{d,\lambda}=0$, it was crucial that $r_{d}$ maps all squared nonzero singular values of $A$ to zero. This is guaranteed by Assumption (Y). Assuming only $Y\neq 0$ and defining the maximal iteration index as in Blanchard and Krämer [5, Section 5.2] would not suffice.
For later purposes we also need the following relatively rough bound. The proof of Lemma 5.5 can be found in Section 9.4.
Lemma 5.5.
We have $\lVert\mathsf{r_{t,<}}g\rVert^{2}\leqslant 6S_{t,\lambda}+2A_{t,\lambda}$ for all $t\in[0,d]$.
5.2 An oracle minimax bound for the prediction error
In this section, we analyse the prediction error of the balanced oracle $\tau_{\mathfrak{b}}$ from Definition 3.3.
Lemma 5.6.
The balanced oracle satisfies without further assumptions
$\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\leqslant\tfrac{4e}{e-1}\inf_{\rho>0}\big{(}\lVert\mathsf{\exp(-\rho x/2)}g\rVert^{2}\vee\lVert\mathsf{(1-\exp(-\rho x))^{1/2}}\xi\rVert^{2}\big{)}.$
Proof.
[…]
$\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}$ $\leqslant 2(A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda})$
$=4\big{(}A_{\tau_{\mathfrak{b}},\lambda}\wedge S_{\tau_{\mathfrak{b}},\lambda}\big{)}$
$\leqslant 4\big{(}\lVert\mathsf{\exp(-\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert x/2)}g\rVert^{2}\wedge\lVert(\mathsf{(\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}\big{)}$ (5.1)
$\leqslant 4\inf_{\rho>0}\big{(}\lVert\mathsf{\exp(-\rho x/2)}g\rVert^{2}\vee\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{)}.$
It remains to apply $y\wedge 1\leqslant\frac{e}{e-1}(1-e^{-y})$ for $y=\rho x\geqslant 0$. ∎
Remark 5.7.
Equation (5.1) yields
$\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\leqslant 4\lVert(\mathsf{(\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert x)^{1/2}\wedge\mathsf{1}})\xi\rVert^{2}\leqslant 4\lVert\xi\rVert^{2}.$
In particular, stopping at the balanced oracle does not incur a significantly larger error than iterating until the terminal iteration index $d$, where the data are perfectly reproduced. Under (DN), the deterministic upper bound $4\delta^{2}$ follows.
Theorem 5.8.
Suppose $f$ satisfies the source condition $\lVert f\rVert_{\mu}\leqslant R$. For all $t\in(0,d]$, we have
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}+2\lVert(\mathsf{(\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}.$ (5.2)
Under (PSD($p$,00,$C_{A}$)) with $p>1/2$ and (GN), the balanced oracle satisfies
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 4\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]\leqslant C_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ (5.3)
with a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$.
Proof.
By 5.3(g) and 4.8, the approximation error satisfies
$A_{t,\lambda}\leqslant R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}.$ (5.4)
Combining the bound in 5.3(d) and (5.4) with Proposition 5.2 yields the first claim.
Assume (GN) and (PSD($p$,00,$C_{A}$)) with $p>1/2$. Since the upper bound for $A_{t,\lambda}$ is decreasing in $\lvert r_{t}^{\prime}(0)\rvert$ and the upper bound for $S_{t,\lambda}$ is increasing in $\lvert r_{t}^{\prime}(0)\rvert$ and $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$, we obtain by Proposition 5.2
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}$ $\leqslant 2\operatorname{{\mathbb{E}}}[A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}]$ (5.5)
$=4\operatorname{{\mathbb{E}}}[A_{\tau_{\mathfrak{b}},\lambda}\wedge S_{\tau_{\mathfrak{b}},\lambda}]$
$\leqslant 4\operatorname{{\mathbb{E}}}\big{[}R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert^{-2\mu-1}\wedge\lVert(\mathsf{(\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}$
$\leqslant 4\operatorname{{\mathbb{E}}}\Big{[}\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}\vee\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{)}\Big{]}$
$\leqslant 4\inf_{\rho>0}\operatorname{{\mathbb{E}}}\big{[}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}$
$\leqslant 4\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+C_{A}^{2}\delta^{2}\sum_{i=1}^{D}\big{(}(\rho i^{-2p})\wedge 1\big{)}\Big{)}.$
Note that $\rho x^{-2p}\leqslant 1$ if and only if $x\geqslant\rho^{1/(2p)}$. A Riemann sum approximation yields
[…]
Hence,
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 4\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+C_{A}^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{1/(2p)}\big{)}.$
The sum is minimal for
$\rho=\big{(}2C_{A}^{-2}(2p-1)(\mu+1/2)^{2\mu+2}R^{2}\delta^{-2}\big{)}^{2p/(4\mu p+2p+1)}>0,$
which yields the rate in (5.3). Inserting $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$ in (5.5), the second claim follows. ∎
Remark 5.9.
As we see in the bound (5.2) on the prediction error, the quantity $\lvert r_{t}^{\prime}(0)\rvert$ controls the size of the approximation and the stochastic error term. The choice of the iteration index $t$ therefore reflects a tradeoff between both errors that can be compared to the bias-variance tradeoff. For deterministic inverse problems, Hanke [10] calls $\lvert r_{k}^{\prime}(0)\rvert$, $k\in{\mathbb{N}}$, the modulus of convergence.
Remark 5.10.
Note that the distribution of the noise $\xi$ only needs to satisfy $\operatorname{{\mathbb{E}}}[\xi_{i}^{2}]=\delta^{2}$, $i=1,\dots,D$, to prove (5.3). So, this bound holds for any centred error $\xi$ with covariance matrix $\delta^{2}I_{D}$. In particular, no Gaussian-type high probability bound is needed to establish the error rate, as required in the statistical CG literature before.
Let us also emphasise that the bounds of Theorem 5.8 are dimension-free because we assumed (PSD($p$,00,$C_{A}$)) with $p>1/2$ so that $AA^{\top}$ is trace class and (5.6) holds. This leads to the conjecture that the same result also holds in the infinite-dimensional Hilbert space setting.
We have shown in Theorem 5.8 that the prediction estimator $\widehat{g}_{\tau_{\mathfrak{b}}}=A\widehat{f}_{\tau_{\mathfrak{b}}}$ at the balanced oracle attains the error rate $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$. The following proposition, which is proved in Section 9.4, gives the corresponding lower bound.
Proposition 5.11.
Assume (PSD($p$,$c_{A}$$\infty$)) with $p>0$ and (GN). We have
$\inf_{\widehat{f}}\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A(\widehat{f}-f)\rVert^{2}\big{]}\geqslant c_{\mu,p,c_{A}}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$
for $\delta/R\leqslant d_{\mu,p,c_{A}}$, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$, where the infimum is taken over all ${\mathbb{R}}^{P}$-valued estimators $\widehat{f}$. The positive constants $c_{\mu,p,c_{A}}$, $d_{\mu,p,c_{A}}$ depend only on $\mu$, $p$ and $c_{A}$.
Remark 5.12.
In the proof of Proposition 5.11, we need to consider alternatives in the first
$t_{\mu,p,R}^{\mathrm{pred}}(\delta)\cong_{\mu,p,c_{A}}(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$
SVD entries of the signal such that we assume that $D$ is at least of this order. In this case, the error rate $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ is minimax.
In the case $D\ll t_{\mu,p,R}^{\mathrm{pred}}(\delta)$, we have $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ and thus by Remark 5.7 also $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{\tau_{b}}-f)\rVert^{2}]\lesssim\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$. In fact, the dimension is so small that regularisation is not needed and in a statistical sense the problem becomes parametric.
Remark 5.13 (Comparison with Showalter’s method).
The CG algorithm can be compared to Showalter’s method (or asymptotic regularisation), which is just the classical gradient flow and appears as continuous-time limit of standard iterative linear regularisation schemes. The Showalter estimator of $f$ at $s\geqslant 0$ is given by $\widehat{f}_{s}^{\mathrm{SW}}=A^{\dagger}(\mathsf{1-\exp(-sx)})Y$ [8, Example 4.7]. Its prediction error can be decomposed as
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}\big{]}=\lVert\mathsf{e^{-sx}}Af\rVert^{2}+\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{1-e^{-sx}})\xi\rVert^{2}\big{]}\eqqcolon B_{s,\lambda}^{2}(f)+V_{s,\lambda}.$ (5.7)
Identifying $s\triangleq\tfrac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, the bias term $B_{s,\lambda}^{2}(f)$ coincides with the upper bound on $A_{t,\lambda}$ given in 5.3(g). Using that $1-e^{-y}\leqslant y\wedge 1\leqslant\tfrac{e}{e-1}(1-e^{-y})$ for $y\geqslant 0$, the stochastic error term $\lVert(\mathsf{1-e^{-sx}})\xi\rVert^{2}$ of $\widehat{f}_{s}^{\mathrm{SW}}$ is up to constants equal to $\lVert(\mathsf{sx\wedge 1})\xi\rVert^{2}$. For $s\triangleq\frac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, this is smaller than the upper bound $\lVert(\mathsf{(\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$ on $S_{t,\lambda}$ from 5.3(d).
We achieve, however, the same order of the oracle prediction error under mild conditions on the singular value decay. Assuming distinct singular values $(\lambda_{j})$ and
$\sum_{l=j}^{D}{\lambda}_{l}^{2}\lesssim j\lambda_{j}^{2},\quad j=2,\dots,D,$ (5.8)
we obtain for $\rho^{-1}\in({\lambda}_{j}^{2},{\lambda}_{j-1}^{2}]$
$\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}=\delta^{2}\Big{(}\#\{i=1,\dots,D\,|\,\lambda_{i}^{2}\geqslant\rho^{-1}\}+\sum_{i:\,\lambda_{i}^{2}<\rho^{-1}}\rho\lambda_{i}^{2}\Big{)}$
$\leqslant\delta^{2}\Big{(}(j-1)+{\lambda}_{j}^{-2}\sum_{l=j}^{D}{\lambda}_{l}^{2}\Big{)}\lesssim\delta^{2}(j-1)=\operatorname{{\mathbb{E}}}\big{[}\lVert{\bf 1}(\mathsf{\rho x\geqslant 1})\xi\rVert^{2}\big{]}\leqslant\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)\wedge 1})\xi\rVert^{2}\big{]}.$ (5.9)
Combining all possible intervals for $\rho^{-1}$ and considering the case $\rho^{-1}\leqslant{\lambda}_{D}^{2}$ separately, the inequality (5.9) holds for all $\rho\geqslant\lambda_{1}^{-2}$. If $\tau_{\mathfrak{b}}\geqslant 1$, we have $\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert\geqslant x_{1,1}^{-1}\geqslant\lambda_{1}^{-2}$ by 4.5(g) and we can apply (5.9) for $\rho\triangleq\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert$. We then obtain by the monotonicity argument with respect to $\rho$ in the proof of Theorem 5.8 and the bias-variance decomposition (5.7) that
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 2\operatorname{{\mathbb{E}}}[A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}]\lesssim\inf_{s\geqslant\lambda_{1}^{-2}}\big{(}B_{s,\lambda}^{2}(f)+V_{s,\lambda}\big{)}=\inf_{s\geqslant\lambda_{1}^{-2}}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}\big{]}.$
It is not difficult to see that for sufficiently small $\delta>0$ we will have $\tau_{\mathfrak{b}}\geqslant 1$ and $\operatorname*{argmin}_{s\geqslant 0}\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}]\geqslant\lambda_{1}^{-2}$ whenever the signal $f$ is not trivial in the sense that at least two SVD entries $f_{i}$ are nonzero.
The interesting fact is that the constant involved does not depend on the signal $f$ so that, in particular, all uniform convergence rates over classes of signals for Showalter’s method under (GN) directly transfer to CG, e.g. under generalised source conditions [19].
The singular value condition (5.8) is easily verified under polynomial singular value decay (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, but is also true in the case of exponential or Gaussian-type decay. Specifically, for exponential decay $\lambda_{i}\thicksim e^{-pi}$, $p>0$, and polynomial source condition we obtain the prediction error rate
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\lesssim_{\mu,p}\log(e+R\delta^{-1})\delta^{2}.$
This rate is minimax optimal, the lower bound follows exactly as Proposition 5.11 by replacing the condition on the singular values accordingly, compare also Tsybakov [25, Theorem 1].
Proposition 5.14.
Suppose $f$ satisfies the source condition $\lVert f\rVert_{\mu}\leqslant R$. Under (DN), for all $t\in(0,d]$, we can bound
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}+2\delta^{2}$
and
$R_{t}\leqslant R(\mu+1/2)^{\mu+1/2}\lvert r_{t}^{\prime}(0)\rvert^{-\mu-1/2}+\delta.$ (5.10)
Proof.
The first part immediately follows from (5.2). Proposition 4.6 implies
$R_{t}\leqslant\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert+\lVert\xi\rVert.$
Corollary 4.8 and Assumption (DN) yield the second claim. ∎
Remark 5.15.
In (5.10), we obtain the same upper bound on the residual norm as Engl, Hanke and Neubauer [8, Lemma 7.10] with a slightly smaller constant on the first summand.
6 Early stopping and its prediction error
In Section 5, we showed that the balanced oracle achieves the optimal MSE rate. Similarly as Blanchard, Hoffmann and Reiß [3, 4], where the balanced oracle stopping index was deterministic though, we mimic $\tau_{\mathfrak{b}}$ by the fully data-driven stopping rule $\tau$ from Definition 3.4.
Lemma 6.1.
The balanced oracle $\tau_{\mathfrak{b}}$ can be equally expressed as
$\tau_{\mathfrak{b}}=\inf\big{\{}t\in[0,d]\,|\,R_{t}^{2}\leqslant\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{t,<}}g}\rangle\big{\}}.$
Proof.
By the definition of $A_{t,\lambda}$, for any $t\in[0,d]$, we deduce
$R_{t}^{2}-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{t,<}}g}\rangle$ $=A_{t,\lambda}+\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}-\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{t,<}}g}\rangle$
$=A_{t,\lambda}+\lVert\mathsf{r_{t,<}^{1/2}}\xi\rVert^{2}-\lVert\xi\rVert^{2}$
$=A_{t,\lambda}-S_{t,\lambda}.$ (6.1)
We find back the original definition of $\tau_{\mathfrak{b}}$. ∎
This alternative expression of $\tau_{\mathfrak{b}}$ motivates the choice of the residual-based stopping rule $\tau$, as explained in Remark 6.2. Note that $R_{\tau}^{2}=\kappa$ and $R_{\tau_{\mathfrak{b}}}^{2}=\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle$ hold by the continuity of $R_{t}^{2}$ in $t$.
Remark 6.2.
The integer-valued discrepancy principle $\lceil\tau\rceil$ with $\kappa\coloneqq c^{2}\delta^{2}$, $c>1$, is widely used in deterministic inverse problems, see Engl, Hanke and Neubauer [8, Section 4.3].
In Lemma 6.1, the cross term $\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle$ will be of negligible order, and we would like to set a threshold $\kappa$ close to $\lVert\xi\rVert^{2}$ for the stopping rule $\tau$. In the Gaussian model (GN) we have $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D$ and $\kappa\coloneqq\delta^{2}D$ is a natural choice. In practice, the noise level $\delta^{2}$ needs to be estimated. As pursued in Blanchard, Hoffmann and Reiß [3, 4], we analyse $\tau$ for general $\kappa>0$ and will later derive error bounds involving the deviation $\lvert\kappa-\delta^{2}D\rvert$.
First, let us determine how much the prediction estimators at the stopping rule $\tau$ and at the balanced oracle $\tau_{\mathfrak{b}}$ differ.
Proposition 6.3.
Under (GN), we have
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-\widehat{f}_{\tau_{\mathfrak{b}}})\rVert^{2}\big{]}\leqslant 2\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]+\delta^{2}\sqrt{2D}+\lvert\kappa-\delta^{2}D\rvert.$
Proof.
Lemma 4.3 implies
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-\widehat{f}_{\tau_{\mathfrak{b}}})\rVert^{2}\big{]}$ $\leqslant\operatorname{{\mathbb{E}}}\big{[}\lvert R_{\tau}^{2}-R_{\tau_{\mathfrak{b}}}^{2}\rvert\big{]}$
$=\operatorname{{\mathbb{E}}}\big{[}\lvert\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle-\kappa\rvert\big{]}$
$\leqslant 2\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]+\operatorname{{\mathbb{E}}}\big{[}\lvert\lVert\xi\rVert^{2}-\delta^{2}D\rvert\big{]}+\lvert\kappa-\delta^{2}D\rvert$
$\leqslant 2\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]+\delta^{2}\sqrt{2D}+\lvert\kappa-\delta^{2}D\rvert,$
where we used $\operatorname{{\mathbb{E}}}[\lvert\lVert\xi\rVert^{2}-\delta^{2}D\rvert]\leqslant\operatorname{Var}(\lVert\xi\rVert^{2})^{1/2}=\delta^{2}\sqrt{2D}$ under (GN). ∎
Remark 6.4.
A first idea could be to bound the cross term via Cauchy–Schwarz inequality:
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]\leqslant\operatorname{{\mathbb{E}}}\big{[}\lVert\xi\rVert^{2}\big{]}^{1/2}\operatorname{{\mathbb{E}}}\big{[}\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}\big{]}^{1/2}.$ (6.2)
Due to $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]^{1/2}=\delta\sqrt{D}$, this is unfortunately highly suboptimal under (GN). If $r_{\tau_{\mathfrak{b}},<}$ were independent of $\xi$ or even deterministic as for the cross term in linear methods [3, 4], then the inner product would have expectation zero and standard deviation $\delta\operatorname{{\mathbb{E}}}[\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}]^{1/2}$, a dimensionless and much tighter bound than (6.2). In our case of intricate randomness of $r_{\tau_{\mathfrak{b}},<}$, we are going to show a slightly larger bound that grows like $\sqrt{\log d}$ in the dimension, but still establishes a sufficiently low order of the cross term, see Proposition 6.8. The probabilistic proof uses concentration of self-normalised Gaussian processes by analysing the potential complexity of $\mathsf{r_{\tau_{\mathfrak{b}},<}}g$ in ${\mathbb{R}}^{D}$. Before, a high probability upper bound on $\lvert r_{t}^{\prime}(0)\rvert$ will be established, giving in Corollary 6.6 an at least linear growth of $t\mapsto S_{t,\lambda}$.
In the sequel, we sometimes write $\lambda(\boldsymbol{\cdot})^{-2}$ instead of $\lambda_{\boldsymbol{\cdot}}^{-2}$ for notational convenience.
Lemma 6.5.
Under (GN), we have for all $z>0$ with probability at least $1-e^{-z}$
$\forall\,t\in[0,d]:\;\lvert r_{t}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
Proof.
For $t=0$, $\lvert r_{t}^{\prime}(0)\rvert=0$ holds. Let $t>0$ and fix $z>0$. By definition of $S_{t,\lambda}$ and Lemma 4.7, the stochastic error can be lower bounded as
$S_{t,\lambda}=\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2}\geqslant\lVert\mathsf{(1-\exp(-\lvert r_{t}^{\prime}(0)\rvert x))^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\sum_{i:\,\lambda_{i}^{-2}\leqslant\lvert r_{t}^{\prime}(0)\rvert}\xi_{i}^{2}.$
Let $I_{\rho}\coloneqq\{i=1,\dots,D\,|\,\lambda_{i}^{-2}\leqslant\rho\}$, $N(\rho)\coloneqq\#I_{\rho}$, and introduce the events
$\Omega(\rho)\coloneqq\Big{\{}\sum_{i\in I_{\rho}}(\delta^{-2}\xi_{i}^{2}-1)\geqslant-2\sqrt{N(\rho)(z+\log d)}\Big{\}},\quad\rho>0.$
By the $\chi^{2}$-concentration result of Laurent and Massart [16, Lemma 1], for fixed $\rho\geqslant\lambda_{1}^{-2}$, we obtain
$\forall\,\zeta>0:\;\operatorname{{\mathbb{P}}}\Big{(}\sum_{i\in I_{\rho}}(\delta^{-2}\xi_{i}^{2}-1)\leqslant-2\sqrt{N(\rho)\zeta}\Big{)}\leqslant e^{-\zeta}.$
Since $\Omega(\rho)$ only changes at $\rho=\tilde{\lambda}_{i}^{-2}$, $i=1,\dots,d$, Bonferroni’s inequality yields
$\operatorname{{\mathbb{P}}}\Big{(}\bigcap_{\rho>0}\Omega(\rho)\Big{)}=\operatorname{{\mathbb{P}}}\Big{(}\bigcap_{i=1}^{d}\Omega(\tilde{\lambda}_{i}^{-2})\Big{)}\geqslant 1-\sum_{i=1}^{d}\operatorname{{\mathbb{P}}}\big{(}\Omega(\tilde{\lambda}_{i}^{-2})^{c}\big{)}\geqslant 1-de^{-(z+\log d)}=1-e^{-z}.$
On $\bigcap_{\rho>0}\Omega(\rho)\subset\Omega(\lvert r_{t}^{\prime}(0)\rvert)$, this gives
$\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}\geqslant N(\lvert r_{t}^{\prime}(0)\rvert)-2\sqrt{(z+\log d)N(\lvert r_{t}^{\prime}(0)\rvert)}.$
The inequality
$N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant 2\Big{(}N(\lvert r_{t}^{\prime}(0)\rvert)-2\sqrt{(z+\log d)N(\lvert r_{t}^{\prime}(0)\rvert)}\Big{)}+4(z+\log d),$
derived from the binomial formula, yields further
$N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d).$
In case $\lvert r_{t}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, we additionally have $N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant D-1$ such that by definition $\lvert r_{t}^{\prime}(0)\rvert<\lambda(N(\lvert r_{t}^{\prime}(0)\rvert)+1)^{-2}$. In case $\lvert r_{t}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, we have $N(\lvert r_{t}^{\prime}(0)\rvert)=D$. By monotonicity of $\lambda(\cdot)^{-2}$, we conclude
$\lvert r_{t}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}$
on the event $\bigcap_{\rho>0}\Omega(\rho)$, which has probability at least $1-e^{-z}$. ∎
Corollary 6.6.
Under (GN), we have for all $z>0$ with probability at least $1-e^{-z}$
$k\leqslant\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1,\quad k=1,\dots,d.$
[…]
$k\lesssim_{p,c_{A},C_{A}}(\delta^{-2}S_{k,\lambda})^{2p/(2p+1)}+(\log d)^{2p/(2p+1)}+z^{2p/(2p+1)},\quad k=1,\dots,d.$
Proof.
[…]
$\lvert r_{k}^{\prime}(0)\rvert=\sum_{i=1}^{k}x_{i,k}^{-1}\geqslant\sum_{i=1}^{k}\lambda_{i}^{-2}\geqslant\lambda_{k}^{-2},\quad k=1,\dots,d.$
Lemma 6.5 yields with probability at least $1-e^{-z}$
$\lambda_{k}^{-2}\leqslant\lvert r_{k}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
By the monotonicity of $\lambda(\cdot)^{-2}$, we obtain the first claim.
Assuming additionally (PSD($p$,$c_{A}$,$C_{A}$)) with $p>0$, we have $\lvert r_{k}^{\prime}(0)\rvert\geqslant\sum_{i=1}^{k}\lambda_{i}^{-2}\gtrsim_{p,C_{A}}k^{2p+1}$ such that we can conclude the second claim again by Lemma 6.5 on the same event. ∎
We obtain the following high probability result for the inner product at general random ‘times’ $\sigma$. Its second part will be used for controlling the reconstruction error.
Theorem 6.7.
Consider the Gaussian noise model (GN). Let $\sigma$ be any random variable taking values in $(0,d]$ (possibly depending on $\xi$). Without any assumption on $(\lambda_{i})$, we have for all $z\geqslant 1$ with probability at least $1-2e^{-z}$
$\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert\leqslant C\delta\lVert\mathsf{r_{\sigma,<}}g\rVert\Big{(}\sqrt{\delta^{-2}S_{\sigma,\lambda}}+\sqrt{\log d}+\sqrt{z}\Big{)}$ (6.3)
with a numerical constant $C>0$. Assuming (PSD($p$,$c_{A}$,$C_{A}$)) with $p>0$ and $\lvert r_{\sigma}^{\prime}(0)\rvert<\lambda_{D}^{-2}$ in addition, we have on the same event
$\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert\leqslant C_{p,c_{A},C_{A}}\delta\lVert\mathsf{r_{\sigma,<}}g\rVert\big{(}(\delta^{-2}S_{\sigma,\lambda})^{p/(2p+1)}+\sqrt{\log d}+\sqrt{z}\big{)}$ (6.4)
with a constant $C_{p,c_{A},C_{A}}>0$ depending only on $p$, $c_{A}$ and $C_{A}$.
Proof.
Fix $z\geqslant 1$. Let us consider the space
$Q_{g,k,i}\coloneqq\bigg{\{}\sum_{j=i+1}^{d}p_{k}(\tilde{\lambda}_{j}^{2})\sum_{l=1,\dots,D:\,\lambda_{l}=\tilde{\lambda}_{j}}\langle{g},{u_{l}}\rangle u_{l}\,\bigg{|}\,p_{k}\in\operatorname{Pol}_{k}\bigg{\}}\subset{\mathbb{R}}^{D}$
for $k=1,\dots,d$, $i=0,\dots,d$. Note that each $Q_{g,k,i}$ is an at most $(k+1)$-dimensional subspace, since $\operatorname{Pol}_{k}$ has dimension $k+1$. We define
$i_{\sigma}\coloneqq\sup\{i=1,\dots,d\,|\,\tilde{\lambda}_{i}^{2}\geqslant x_{1,\sigma}\}\vee 0.$
The definition of $r_{\sigma,<}$ implies
$r_{\sigma,<}g=\sum_{j=i_{\sigma}+1}^{d}r_{\sigma}(\tilde{\lambda}_{j}^{2})\sum_{l=1,\dots,D:\,\lambda_{l}=\tilde{\lambda}_{j}}\langle{g},{u_{l}}\rangle u_{l}\in Q_{g,\lceil\sigma\rceil,i_{\sigma}}.$
Denote by $\Pi_{g,k,i}$ the orthogonal projection of ${\mathbb{R}}^{D}$ onto $Q_{g,k,i}$. By writing
$\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert=\lVert\mathsf{r_{\sigma,<}}g\rVert\frac{\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert}{\lVert\mathsf{r_{\sigma,<}}g\rVert}\leqslant\lVert\mathsf{r_{\sigma,<}}g\rVert\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert,$ (6.5)
we can reduce the problem to bounding the norm of the projection $\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert=\delta\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}Z\rVert$. Note that $\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\leqslant\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert^{2}]^{1/2}\leqslant\sqrt{k+1}$ because $\lVert\Pi_{g,k,i}Z\rVert^{2}$ is $\chi^{2}$-distributed with $\dim(Q_{g,k,i})$ degrees of freedom for deterministic $k$ and $i$. Moreover, the function $x\mapsto\lVert\Pi_{g,k,i}x\rVert$, $x\in{\mathbb{R}}^{D}$, is $1$-Lipschitz such that we obtain by Gaussian concentration [18, Theorem 3.4] and Bonferroni’s inequality
[…]
$\leqslant\sum_{k=1}^{d}\sum_{i=0}^{d}\operatorname{{\mathbb{P}}}\Big{(}\lVert\Pi_{g,k,i}Z\rVert-\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\geqslant\big{(}2z+2\log(d(d+1))\big{)}^{1/2}\Big{)}$
$\leqslant d(d+1)e^{-(2z+2\log(d(d+1)))/2}=e^{-z}.$
Hence, with probability at least $1-e^{-z}$
$\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert$ $\leqslant\delta\Big{(}\sqrt{\lceil\sigma\rceil+1}+\big{(}2z+2\log(d(d+1))\big{)}^{1/2}\Big{)}$
$\lesssim\delta\big{(}\sqrt{\sigma}+\sqrt{\log d}+\sqrt{z}\big{)}$ (6.6)
holds. Without any assumption on $(\lambda_{i})$, we have by Corollary 6.6
$\sigma\leqslant\lfloor\sigma\rfloor+1\lesssim\delta^{-2}S_{\sigma,\lambda}+\log d+z$ (6.7)
also with probability at least $1-e^{-z}$, where we used that $t\mapsto S_{t,\lambda}$ is nondecreasing by 3.2(c). Combining (6.5)–(6.7) yields the first claim. For the second assertion, we use $\lvert r_{\lfloor\sigma\rfloor}^{\prime}(0)\rvert\leqslant\lvert r_{\sigma}^{\prime}(0)\rvert<\lambda_{D}^{-2}$ by assumption and obtain by Corollary 6.6 under (PSD($p$,$c_{A}$,$C_{A}$))
$\sigma\leqslant\lfloor\sigma\rfloor+1\lesssim_{p,c_{A},C_{A}}(\delta^{-2}S_{\sigma,\lambda})^{2p/(2p+1)}+\log d+z$ (6.8)
on the same event where (6.7) holds. By (6.5), (6.6) and (6.8), the claim is deduced. ∎
We conclude the required bound in expectation for the cross term.
Proposition 6.8.
For the Gaussian noise model (GN), without any assumption on $(\lambda_{i})$, the bound
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]\leqslant C\delta^{2}\big{(}\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]+\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]^{1/2}\sqrt{\log d}\big{)}$
holds true with a numerical constant $C>0$.
Proof.
From Lemma 5.5 and the identity $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$, we deduce
$\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}\leqslant 8S_{\tau_{\mathfrak{b}},\lambda}.$
We obtain by (6.3) and Lemma 9.7
$\operatorname{{\mathbb{E}}}\bigg{[}\frac{\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert^{2}}{\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}}\bigg{]}\lesssim\delta^{2}\big{(}\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]+\log d\big{)}.$
By the Cauchy–Schwarz inequality, we conclude
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]$ $\leqslant\operatorname{{\mathbb{E}}}\big{[}\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}\big{]}^{1/2}\operatorname{{\mathbb{E}}}\bigg{[}\frac{\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert^{2}}{\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}}\bigg{]}^{1/2}$
$\lesssim\delta^{2}\big{(}\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]+\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]^{1/2}\sqrt{\log d}\big{)}.\qed$
Finally, we can derive the oracle bound for the prediction error.
Theorem 6.9.
For the Gaussian noise model (GN), without any assumption on $(\lambda_{i})$, the prediction error at the stopping rule $\tau$ satisfies the balanced oracle inequality
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\big{]}\leqslant C\big{(}\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]+\delta^{2}\sqrt{D}+\lvert\kappa-\delta^{2}D\rvert\big{)}$
with a numerical constant $C>0$.
Proof.
Propositions 6.3 and 6.8 give
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-\widehat{f}_{\tau_{\mathfrak{b}}})\rVert^{2}\big{]}\lesssim\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]+\delta^{2}\sqrt{D}+\lvert\kappa-\delta^{2}D\rvert.$
By Proposition 5.2 and the identity $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$, we have
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 4\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}].$
Combining both upper bounds with the inequality
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\big{]}\leqslant 2\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-\widehat{f}_{\tau_{\mathfrak{b}}})\rVert^{2}\big{]}+2\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}$
yields the claim. ∎
If $\lvert\kappa-\delta^{2}D\rvert$ is at most of the order of $\delta^{2}\sqrt{D}$, the balanced oracle inequality immediately implies minimax adaptivity of the early stopping rule for the prediction error over Sobolev-type ellipsoids for a range of regularity parameters $\mu$.
Corollary 6.10.
Assume (PSD($p$,00,$C_{A}$)) with $p>1/2$, (GN) and $\lvert\kappa-\delta^{2}D\rvert\leqslant C_{\kappa}\delta^{2}\sqrt{D}$ for some $C_{\kappa}\geqslant 0$. Under the source condition $\lVert f\rVert_{\mu}\leqslant R$, we have
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\big{]}\leqslant C_{\mu,p,C_{A},C_{\kappa}}\big{(}\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)+\delta^{2}\sqrt{D}\big{)}$
with a constant $C_{\mu,p,C_{A},C_{\kappa}}>0$ depending only on $\mu$, $p$, $C_{A}$ and $C_{\kappa}$. Under (PSD($p$,$c_{A}$,$C_{A}$)), this gives the minimax optimal rate for all regularities ${\mu>0}$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
Proof.
Combining the bound on $\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]$ from Theorem 5.8 with Theorem 6.9 yields the upper bound. If $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, then $\delta^{2}\sqrt{D}\lesssim\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$, and we obtain minimax adaptivity for the given range of regularity parameters $\mu$ by Proposition 5.11. ∎
Proposition 6.11.
Under (DN), the discrepancy principle $\tau$ with $\kappa\coloneqq c^{2}\delta^{2}$, $c\geqslant 0$, satisfies
$\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}\leqslant(c+1)^{2}\delta^{2}.$
Proof.
The definition of $\tau$ and the assumption $\lVert\xi\rVert\leqslant\delta$ directly yield
$\lVert A(\widehat{f}_{\tau}-f)\rVert=\lVert A\widehat{f}_{\tau}-Y+\xi\rVert\leqslant R_{\tau}+\lVert\xi\rVert\leqslant(c+1)\delta.\qed$
7 Transfer to the reconstruction error
7.1 Bound for the reconstruction error
The main tool to control the reconstruction error is the following relatively rough bound in terms of an approximation error for $f^{\dagger}$ and the quantities already studied for the prediction error.
Proposition 7.1.
For $t\in[0,d]$, we have
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}\leqslant 2\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}+4\lvert r_{t}^{\prime}(0)\rvert S_{t,\lambda}+2\lvert r_{t}^{\prime}(0)\rvert A_{t,\lambda}.$
Proof.
For $t=0$, the claim is clear. Let $t>0$. Since $1-r_{t}$ is a polynomial vanishing in zero, we have
$A^{\dagger}(1-r_{t})(AA^{\top})A=A^{\top}\Big{(}\tfrac{1-r_{t}(x)}{x}\Big{)}(AA^{\top})A=(1-r_{t})(A^{\top}A)$
and thus
$\widehat{f}_{t}-f^{\dagger}=A^{\dagger}(1-r_{t})(AA^{\top})(Af^{\dagger}+\xi)-f^{\dagger}=-r_{t}(A^{\top}A)f^{\dagger}+A^{\dagger}(1-r_{t})(AA^{\top})\xi.$
We conclude by direct calculation
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}$
$=\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{<}^{2}+\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{>}^{2}$
$\leqslant\lVert r_{t}(A^{\top}A)f^{\dagger}-A^{\dagger}(I_{D}-r_{t}(AA^{\top}))\xi\rVert^{2}_{<}+x_{1,t}^{-1}\lVert A(\widehat{f}_{t}-f^{\dagger})\rVert^{2}$
$\leqslant 2\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}+2\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))_{<}\xi\rVert^{2}+\lvert r_{t}^{\prime}(0)\rvert\lVert A(\widehat{f}_{t}-f)\rVert^{2},$ (7.1)
where we used for the last inequality that $\lvert r_{t}^{\prime}(0)\rvert\geqslant x_{1,t}^{-1}$. Since additionally $x^{-1}(1-r_{t}(x))_{<}\leqslant\lvert r_{t}^{\prime}(0)\rvert$ by Lemma 4.7, the second term satisfies
$\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))_{<}\xi\rVert^{2}=\lVert\mathsf{x^{-1/2}(1-r_{t})_{<}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert S_{t,\lambda}.$ (7.2)
Proposition 5.2 yields for the last term
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2(A_{t,\lambda}+S_{t,\lambda}).$
It remains to combine these inequalities. ∎
Remark 7.2.
The approach to bound the reconstruction error as in (7.1) is similar to that in the proof of Engl, Hanke and Neubauer [8, Lemma 7.11]. Yet, instead of splitting the norm at $x_{1,t}$, they choose a positive $\varepsilon\leqslant\lvert r_{t}^{\prime}(0)\rvert^{-1}$ such that their upper bound is minimal. This choice is necessary for their approach since they use, under the source condition $\lVert f\rVert_{\mu}\leqslant R$, the rough bound
$\lVert(\mathsf{r_{t}(x)}\mathbf{1}(\mathsf{x}\leqslant\varepsilon))(A^{\top}A)f\rVert\leqslant\sup_{x\in[0,\varepsilon]}\lvert r_{t}(x)x^{\mu}\rvert R\leqslant\varepsilon^{\mu}R.$
Instead, we will bound the term in (7.4) below with respect to the error terms. For the second summand in (7.1), they consider the worst-case bound
$\lVert\mathbf{1}(\mathsf{x\leqslant\varepsilon})\mathsf{x^{-1/2}(1-r_{t}(x))}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert\lVert\xi\rVert^{2},$
which is suboptimal in the stochastic noise setting (GN).
Using an interpolation inequality similar to [8, (2.49)], we find the following bound.
Corollary 7.3.
Assume the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$ for the minimum-norm solution $f^{\dagger}$. Then, for $t\in[0,d]$, we have
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}\leqslant 16R^{2/(2\mu+1)}(S_{t,\lambda}\vee A_{t,\lambda})^{2\mu/(2\mu+1)}+6\lvert r_{t}^{\prime}(0)\rvert(S_{t,\lambda}\vee A_{t,\lambda}).$
Proof.
By Proposition 7.1, it remains to bound $\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}$. Note that $\langle{f^{\dagger}},{v_{i}}\rangle=f_{i}$ for $i=1,\dots,D$ and $\langle{f^{\dagger}},{v_{i}}\rangle=0$ for $i=D+1,\dots,P$ in case of $P>D$. Using the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$ and Hölder’s inequality with the conjugate exponents $\tilde{p}\coloneqq(2\mu+1)/(2\mu)$ and $\tilde{q}\coloneqq 2\mu+1$, we obtain
$\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}$ $=\sum_{i=1}^{D}r_{t,<}^{2}(\lambda_{i}^{2})\lambda_{i}^{4\mu/(2\mu+1)}\lambda_{i}^{-4\mu/(2\mu+1)}f_{i}^{2}$
$\leqslant\Big{(}\sum_{i=1}^{D}r_{t,<}^{2}(\lambda_{i}^{2})\lambda_{i}^{2}f_{i}^{2}\Big{)}^{2\mu/(2\mu+1)}\Big{(}\sum_{i=1}^{D}r_{t,<}^{2}(\lambda_{i}^{2})\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/(2\mu+1)}$
$\leqslant\lVert\mathsf{r_{t,<}}g\rVert^{4\mu/(2\mu+1)}\Big{(}\sum_{i=1}^{D}\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/(2\mu+1)}$
$\leqslant\lVert\mathsf{r_{t,<}}g\rVert^{4\mu/(2\mu+1)}R^{2/(2\mu+1)},$ (7.3)
where in the second last step $\lvert r_{t,<}(x)\rvert\leqslant 1$ was used. Invoking Lemma 5.5, we thus have
$2\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}\leqslant 16R^{2/(2\mu+1)}(S_{t,\lambda}\vee A_{t,\lambda})^{2\mu/(2\mu+1)}$ (7.4)
and conclude the claim. ∎
Corollary 7.4.
Under (DN) and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we have for $t\in(0,d]$
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}\leqslant C_{\mu}\big{(}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu}R^{2}+\lvert r_{t}^{\prime}(0)\rvert\delta^{2}\big{)}$ (7.5)
with a constant $C_{\mu}>0$ depending only on $\mu$.
Proof.
Combining Proposition 7.1, (7.3) and $S_{t,\lambda}\leqslant\delta^{2}$ under (DN), we obtain
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}\lesssim\lVert\mathsf{r_{t,<}}g\rVert^{4\mu/(2\mu+1)}R^{2/(2\mu+1)}+\lvert r_{t}^{\prime}(0)\rvert\delta^{2}+\lvert r_{t}^{\prime}(0)\rvert A_{t,\lambda}.$
Noting $\lVert\mathsf{r_{t,<}}g\rVert\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert$ due to $\lvert r_{t,<}(x)\rvert\leqslant 1$, we conclude by Corollary 4.8 and (5.4)
$\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}$ $\lesssim_{\mu}\big{(}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}R^{2}\big{)}^{2\mu/(2\mu+1)}R^{2/(2\mu+1)}+\lvert r_{t}^{\prime}(0)\rvert\delta^{2}+\lvert r_{t}^{\prime}(0)\rvert^{-2\mu}R^{2}$
$\lesssim_{\mu}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu}R^{2}+\lvert r_{t}^{\prime}(0)\rvert\delta^{2}.\qed$
Remark 7.5.
For small $t$, the first summand in the bound (7.5) dominates and the iterates seem to converge until both summands are of the same order. From then on, the estimator diverges. This phenomenon is named semiconvergence [10, p. 45]. It corresponds to the tradeoff of approximation and stochastic error for the prediction error, see Remark 5.9.
We are in a position to prove convergence rates for the reconstruction error under Gaussian noise.
Theorem 7.6.
Assume (PSD($p$,00,$C_{A}$)) with $p>1/2$ and (GN). Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the reconstruction error at the balanced oracle satisfies
$\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\leqslant C_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$
with a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$. In particular, under (PSD($p$,$c_{A}$,$C_{A}$)), the balanced oracle is simultaneously minimax adaptive for prediction and reconstruction error for all $\mu>0$ with $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
Proof.
Corollary 7.3 and the identity $S_{\tau_{\mathfrak{b}},\lambda}=A_{\tau_{\mathfrak{b}},\lambda}$ yield
$\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\leqslant 16R^{2/(2\mu+1)}\operatorname{{\mathbb{E}}}[S_{\tau_{\mathfrak{b}},\lambda}]^{2\mu/(2\mu+1)}+6\operatorname{{\mathbb{E}}}[\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert S_{\tau_{\mathfrak{b}},\lambda}],$
where we have applied Jensen’s inequality. The first summand has the desired rate by Theorem 5.8. Using 5.3(d), (5.4) and (5.6), we deduce similarly to the proof of Theorem 5.8
$\operatorname{{\mathbb{E}}}[\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert S_{\tau_{\mathfrak{b}},\lambda}]$
$=\operatorname{{\mathbb{E}}}[\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert A_{\tau_{\mathfrak{b}},\lambda}\wedge\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert S_{\tau_{\mathfrak{b}},\lambda}]$
$\leqslant\operatorname{{\mathbb{E}}}\big{[}R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert^{-2\mu}\wedge\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert\lVert(\mathsf{(\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}$
$\leqslant\operatorname{{\mathbb{E}}}\Big{[}\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu}\vee\rho\lVert(\mathsf{(\rho\mathsf{x})^{1/2}\wedge\mathsf{1}})\xi\rVert^{2}\big{)}\Big{]}$
$\leqslant\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu}+C_{A}^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{(2p+1)/(2p)}\big{)}$
$\lesssim_{\mu,p,C_{A}}R^{(4p+2)/(4\mu p+2p+1)}\delta^{8\mu p/(4\mu p+2p+1)},$
which gives the right rate for the second summand. Under (PSD($p$,$c_{A}$,$C_{A}$)), this rate is minimax optimal for the given range of regularity parameters $\mu$, see Remark 7.7. The minimax optimality for the prediction error was shown in Theorems 5.8 and 5.11. ∎
Remark 7.7.
The error rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ coincides with the minimax rate in the infinite-dimensional Gaussian sequence space model, and the corresponding lower bound can be proven in the same way [see 15, Proposition 4.23], very similar to the proof of Proposition 5.11: under (PSD($p$,$c_{A}$$\infty$)) with $p>0$ and (GN), the rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ is minimax optimal, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$. If $D\ll(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, then $\delta^{2}D^{2p+1}\ll\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$, and we obtain better error bounds without regularisation exactly as in Remark 5.12:
$\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\leqslant\lambda_{D}^{-2}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger})\rVert^{2}\big{]}\lesssim_{c_{A}}\delta^{2}D^{2p+1}\ll\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta).$
Also Remark 5.10 applies again for the upper bound: only conditions on the second moments of the $\xi_{i}$ are used and the bound is dimension-free.
7.2 Early stopping and its reconstruction error
We analyse the reconstruction error of the early stopping rule $\tau$ from Definition 3.4. Unfortunately, there is no reconstruction error analogue of the oracle inequality in Theorem 6.9. Instead, we need to directly bound $\lvert r_{\tau}^{\prime}(0)\rvert$ and the errors at $\tau$ in high probability to apply Corollary 7.3. For this purpose Lemmas 6.5 and 6.7 are essential, requiring $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$. The case $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$ leads to high stochastic error and is dealt with separately, assuming the same lower bound on the dimension $D$ as for the lower bound on the reconstruction error in Remark 7.7. The proofs for this subsection are delegated to Section 9.5.
Proposition 7.8.
Assume (GN), (PSD($p$,$c_{A}$,$C_{A}$)) and put $M_{\tau,\lambda}\coloneqq A_{\tau,\lambda}\vee S_{\tau,\lambda}$. Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the following holds:
  1. (a)
    We have for $p>0$ and all $z>0$ with probability at least $1-e^{-z}$
    $\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\lesssim_{p,c_{A}}R^{2/(2\mu+1)}M_{\tau,\lambda}^{2\mu/(2\mu+1)}+\delta^{-4p}S_{\tau,\lambda}^{2p}M_{\tau,\lambda}+(\log d)^{2p}M_{\tau,\lambda}+z^{2p}M_{\tau,\lambda}.$
  2. (b)
    We have for $p>1/2$ and all $z>0$ with probability at least $1-e^{-z}$
    $M_{\tau,\lambda}\lesssim_{\mu,p,C_{A}}R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}+\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert+\delta^{2}z.$
  3. (c)
    We have for $p>1/2$, $\lvert\kappa-\delta^{2}D\rvert\leqslant C_{\kappa}\delta^{2}\sqrt{D}$ for some $C_{\kappa}\geqslant 0$ and all $z\geqslant 1$ with probability at least $1-5e^{-z}$
    $M_{\tau,\lambda}\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}+\delta^{2}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)},$
    provided $D\gtrsim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$.
Remark 7.9.
The argument for 7.8(c) hinges on the fact that the bound on $\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$ derived in (9.9) is of smaller order than the bound on $M_{\tau,\lambda}$ in 7.8(b). This is provided by the bound (6.4), which is valid under (PSD($p$,$c_{A}$,$C_{A}$)), but not under exponential singular value decay.
Inserting the high probability bound of 7.8(c) into 7.8(a) yields a high probability bound for the reconstruction error that in turn implies the main result on the reconstruction error under early stopping.
Theorem 7.10.
Assume (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, (GN) and $\lvert\kappa-\delta^{2}D\rvert\leqslant C_{\kappa}\delta^{2}\sqrt{D}$ for some $C_{\kappa}\geqslant 0$. Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we have for all $z\geqslant 1$ with probability at least $1-6e^{-z}$
$\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\leqslant C_{\mu,p,c_{A},C_{A},C_{\kappa}}\Big{(}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)+\delta^{2}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2p+1}\Big{)}$
and the bound in expectation
$\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\big{]}\leqslant\widetilde{C}_{\mu,p,c_{A},C_{A},C_{\kappa}}\big{(}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)+\delta^{2}D^{p+1/2}\big{)},$
provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim D$. Here $C_{\mu,p,c_{A},C_{A},C_{\kappa}},\widetilde{C}_{\mu,p,c_{A},C_{A},C_{\kappa}}$ are positive constants depending only on $\mu$, $p$, $c_{A}$, $C_{A}$ and $C_{\kappa}$. This gives the minimax rate for all regularities $\mu>0$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
Remark 7.11.
Comparing Theorem 7.10 and Corollary 6.10, we note that the early stopping rule $\tau$ is minimax adaptive both for prediction and reconstruction error over the same range of regularity parameters $\mu$.
The error term $\delta^{2}D^{p+1/2}$ in Theorem 7.10 is due to the difference $\lvert\lVert\xi\rVert^{2}-\delta^{2}D\rvert$ and also appears in other early stopping regularisation methods [3, 23].
Proposition 7.12.
Under (DN) and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the discrepancy principle $\tau$ with $\kappa\coloneqq c^{2}\delta^{2}$, $c>1$, satisfies
$\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\leqslant C_{\mu,c}R^{2/(2\mu+1)}\delta^{4\mu/(2\mu+1)},$
where $C_{\mu,c}>0$ is a constant depending only on $\mu$ and $c$.
Remark 7.13.
Engl, Hanke and Neubauer [8, Theorem 7.12] obtain the same result for noninterpolated CG, yet arguing sometimes differently. In addition, they have to control the difference between two iterates using the update polynomials, which we circumvent by interpolation.
8 Numerical examples
[…]
The SVD representations of the three signals from [3, 4, 23] are given by
$f_{i}^{(1)}$ $=5\exp(-0.1i),$ (supersmooth)
$f_{i}^{(2)}$ $=5000\lvert\sin(0.01i)\rvert i^{-1.6},$ (smooth)
$f_{i}^{(3)}$ $=250\lvert\sin(0.002i)\rvert i^{-0.8},\quad i=1,\dots,D.$ (rough)
The names allude to the decay of the coefficients, interpreted as Fourier coefficients. For every signal we run $1000$ Monte-Carlo simulations.
In addition, we consider the test problem gravity from the popular Matlab package Regularization Tools [11, 12]. The one-dimensional gravity surveying problem leads to the Fredholm integral equation [12, (2.1)] on $L^{2}([0,1])$
[…]
which is discretised by the midpoint quadrature rule. The ill-posedness is controlled by the parameter $\mathtt{d}$. We choose the default value $\mathtt{d}=0.25$. The true signal (mass distribution) is given by
$f^{(4)}(t)=\sin(\pi t)+0.5\sin(2\pi t),\quad t\in[0,1].$
We obtain discretisations $A\in{\mathbb{R}}^{D\times D}$ and $f^{(4)}\in{\mathbb{R}}^{D}$ of the forward operator and the signal, respectively, for dimension $D=4096$. Under (GN) with noise level $\delta=0.01$, we run $1000$ Monte-Carlo simulations. Jahn [14] considers the same example for the early stopped Landweber (classical gradient descent) algorithm.
Figure 1: Relative efficiencies of early stopping for prediction and reconstruction error (higher is better)
In both settings, we calculate the conjugate gradient estimator $\widehat{f}_{\tau}$, where the threshold for the early stopping rule is chosen as $\kappa\coloneqq\delta^{2}D$ for the first three examples and $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$ for the gravity problem. At the latest, the algorithm is terminated at step $T\in{\mathbb{N}}$ if $T$ reaches the maximal iteration index $D$ or if the squared norm $\lVert A^{\top}(Y-A\widehat{f}_{T})\rVert^{2}$ of the transformed residual vector is smaller than $10^{-8}$ (emergency stop). Otherwise, it would break down with division by zero [8, p. 179]. Note that for ill-conditioned matrices, $R_{t}^{2}$ does not have to be close to zero for $\lVert A^{\top}(Y-A\widehat{f}_{t})\rVert^{2}$ to be small.
We measure how close the estimator $\widehat{f}_{\tau}$ is to the optimal iterate along the (continuous) interpolation path $(\widehat{f}_{t},t\in[0,T])$ by calculating the relative efficiencies
$\frac{\min_{t\in[0,T]}\lVert A(\widehat{f}_{t}-f)\rVert}{\lVert A(\widehat{f}_{\tau}-f)\rVert}\quad\text{and}\quad\frac{\min_{t\in[0,T]}\lVert\widehat{f}_{t}-f\rVert}{\lVert\widehat{f}_{\tau}-f\rVert}$
for the prediction and reconstruction error, respectively. Ideally, the values should be close to one.
First, consider the performance for the signals $f^{(1)},f^{(2)},f^{(3)}$ in Figure 1. The relative efficiencies are surprisingly high, especially for the reconstruction error. Note that a relative efficiency larger than $0.5$ means that the error is less than twice the optimal error. For both criteria the supersmooth signal suffers from higher variability than the other two due to its very small optimal errors, compare Table 1. Blanchard, Hoffmann and Reiß [4] analyse the same stopping rule for the Landweber algorithm. Their results for the three signals are very similar, conjugate gradients even perform slightly better.
The relative efficiencies for the gravity example show more variability. Note that this inverse problem is severely ill-posed with exponentially decaying singular values violating our assumptions for the adaptivity in reconstruction error of early stopping. Additionally, the optimal oracle errors are even smaller than for the supersmooth signal $f^{(1)}$, see Table 1. So, the gravity test problem serves as an example where the theory lets us expect bad performance. Further problems arise from machine precision: although in theory the residuals reach zero at iteration $D$ at the latest, in our calculations the residuals tend to stagnate at some nonzero value, which might be larger than the usual threshold $\delta^{2}D$. In these cases, the algorithm is terminated via the aforementioned emergency stop. In unreported simulations with $\kappa\coloneqq\delta^{2}D$, this happened in $44$ percent of the runs.
Remark 7.9 explains that we cannot expect $\lvert S_{\tau,\lambda}-A_{\tau,\lambda}\rvert$ to be small for severely ill-posed problems so that we loose control of $\lvert r_{\tau}^{\prime}(0)\rvert$. Therefore, we recommend to stop slightly earlier and we choose here $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$, still satisfying our condition on $\kappa$ in Corollary 6.10 with $C_{\kappa}=1$, leading to emergency stops in $21$ percent of the runs. If we ignore these runs, the minimal relative efficiency for the reconstruction error is about $0.2$ such that in the worst case the reconstruction error is five times the optimal error. The interesting conclusion is that the early stopping criterion provides reasonable results even in a setting not covered by the theory.
Table 1: Medians (mean absolute deviations) of stopping indices and errors for different signals
Stopping indices Corresponding errors
$t_{\mathrm{e}}^{\mathrm{pred}}$ $t_{\mathrm{e}}^{\mathrm{rec}}$ $\tau$ $\lVert A(\widehat{f}_{t_{\mathrm{e}}^{\mathrm{pred}}}-f)\rVert$ $\lVert A(\widehat{f}_{\tau}-f)\rVert$ $\lVert\widehat{f}_{t_{\mathrm{e}}^{\mathrm{rec}}}-f\rVert$ $\lVert\widehat{f}_{\tau}-f\rVert$
$f^{(1)}$ $6.13$ ($0.33$) $5.57$ ($0.04$) $5.07$ ($0.24$) $0.11$ ($0.01$) $0.18$ ($0.02$) $0.81$ ($0.02$) $0.87$ ($0.05$)
$f^{(2)}$ $15.42$ ($0.64$) $12.63$ ($0.02$) $11.45$ ($0.07$) $0.32$ ($0.01$) $0.41$ ($0.01$) $5.66$ ($0.06$) $6.01$ ($0.08$)
$f^{(3)}$ $19.49$ ($0.19$) $18.16$ ($0.1$) $14.15$ ($0.05$) $0.59$ ($0.01$) $0.67$ ($0.01$) $21.77$ ($0.11$) $22.69$ ($0.12$)
$f^{(4)}$ $10.68$ ($2.14$) $12.43$ ($1.06$) $6.97$ ($2.9$) $0.04$ ($0.02$) $0.09$ ($0.03$) $0.63$ ($0.12$) $1.18$ ($0.29$)
In Table 1, we state the medians (and mean absolute deviations around the median) of the optimal (noise-dependent) stopping indices
$t_{\mathrm{e}}^{\mathrm{pred}}=\operatorname*{argmin}_{t\in[0,T]}\lVert A(\widehat{f}_{t}-f)\rVert\quad\text{and}\quad t_{\mathrm{e}}^{\mathrm{rec}}=\operatorname*{argmin}_{t\in[0,T]}\lVert\widehat{f}_{t}-f\rVert$
and of the stopping rules $\tau$ as well as of the corresponding root mean squared errors for the Monte-Carlo iterations. For all signals we tend to stop rather too early than too late compared to $t_{\mathrm{e}}^{\mathrm{pred}}$ and $t_{\mathrm{e}}^{\mathrm{rec}}$. Blanchard, Hoffmann and Reiß [4] argue that in case of the Landweber method for the first three signals a slightly smaller value of $\kappa$ is appropriate.
Finally, we analyse the behaviour of the errors for $f^{(1)},f^{(2)},f^{(3)}$ in the asymptotic regime where the dimension $D(\delta)$ grows as the noise level $\delta$ vanishes. Following the asymptotical setting of Stankewitz [23], we choose the dimensions $D=D_{m}=100\cdot 2^{m}$ for $m=0,\dots,10$ with noise levels
$\delta_{m}=RD_{m}^{-2\mu p-p-1/2}\quad\text{such that}\quad D_{m}=(R^{2}\delta_{m}^{-2})^{1/(4\mu p+2p+1)},\quad m=0,\dots,10,$
[…]
Figure 2: Log-log plots of the prediction and reconstruction error rates for $f^{(1)}$, $f^{(2)}$ and $f^{(3)}$ (solid lines: risk estimates at $\tau$, dashed lines: risk estimates at $t_{\mathrm{e}}^{\mathrm{pred}}$ and $t_{\mathrm{e}}^{\mathrm{rec}}$, respectively)
For all three signals and both error criteria, the early stopped estimators closely match the optimal ones. For the rough signal, the asymptotic regime applies only for small $\delta$, that is large $D$. This is explained by the fact that the signals still change with $D$, having more nonzero coefficients and thus inducing larger approximation errors. Again, the similarity is even higher for the reconstruction error. The slopes of the minimax rates are well matched, even exhibiting a small asymptotic constant.
9 Remaining results and proofs
For the convenience of the reader, the next two subsections collect and sometimes prove classical results for the CG algorithm.
9.1 Results for zeros of orthogonal polynomials
We consider the Wronskian
$W(x;p,q)\coloneqq p(x)q^{\prime}(x)-p^{\prime}(x)q(x),\quad x\in{\mathbb{R}},$ (9.1)
of two polynomials $p$ and $q$ on ${\mathbb{R}}$. Note that the Wronskian $W(x;p,q)$ is continuous in $x$. The following result is due to Beardon and Driver [1, Section 4].
Lemma 9.1 (Properties of the Wronskian).
If $W(x;p,q)$ is nonzero in an open interval $J\subset{\mathbb{R}}$, then
  1. (a)
    $p$ and $q$ have no common zeros in $J$,
  2. (b)
    any zero of $p$ and $q$ in $J$ is a simple zero,
  3. (c)
    any two consecutive zeros of $p$ in $J$ are separated by a zero of $q$,
  4. (d)
    any two consecutive zeros of $q$ in $J$ are separated by a zero of $p$.
Let $\mu$ denote a discrete measure on the interval $[a,b]\subset{\mathbb{R}}$ with a support consisting of $N<\infty$ points. Denote by $\lVert\cdot\rVert_{L^{2}(\mu)}$ the $L^{2}$-norm with respect to $\mu$.
Lemma 9.2.
Let $(p_{n})_{n=0}^{N}$ be a sequence of orthogonal polynomials in $L^{2}([a,b],\mu)$ with $\lVert p_{n}\rVert_{L^{2}(\mu)}>0$ for $n=0,\dots,N-1$. Then the Wronskian $W(x;p_{n},p_{n+1})$, $n=0,\dots,N-1$, of two consecutive orthogonal polynomials does not vanish on $(a,b)$.
Proof.
Denote by $k_{m}$, $m=0,\dots,N$, the highest coefficient of $p_{m}$. Since $\lVert p_{n}\rVert_{L^{2}(\mu)}>0$, $n=0,\dots,N-1$, we obtain by the Christoffel–Darboux formula [24, Theorem 3.2.2] that
$W(x;p_{n},p_{n+1})$ $=p_{n}(x)p_{n+1}^{\prime}(x)-p_{n}^{\prime}(x)p_{n+1}(x)$
$=\frac{\lVert p_{n}\rVert_{L^{2}(\mu)}^{2}k_{n+1}}{k_{n}}\bigg{(}\sum_{i=1}^{n}\frac{p_{i}^{2}(x)}{\lVert p_{i}\rVert_{L^{2}(\mu)}^{2}}+\frac{k_{0}^{2}}{\lVert p_{0}\rVert_{L^{2}(\mu)}^{2}}\bigg{)}\neq 0.\qed$
9.2 Results for classical conjugate gradients
Recall Definition 3.1 of the noninterpolated estimators. For $k<d$, the minimising polynomial $r_{k}$ is given by the orthogonal projection of the origin on the affine subspace $\operatorname{Pol}_{k,1}$ of $\operatorname{Pol}_{k}$ and is therefore unique. In the case $k=d$, the zeros of $r_{d}$ are given by the distinct squared nonzero singular values of $A$ as shown in 9.5(a). Since $r_{d}(0)=1$, $r_{d}$ is also uniquely determined.
Lemma 9.3.
The residual polynomial $r_{k}$ is orthogonal to $\operatorname{Pol}_{k,0}$ with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$ for $k=0,\dots,d$. We write $r_{k}\perp_{Y}\operatorname{Pol}_{k,0}.$
Proof.
First, consider the case $k<d$. By the definition of $r_{k}$, the polynomial $r_{k}-1$ is the orthogonal projection of $-1$ with respect to $\lVert\cdot\rVert_{Y}$ onto the linear subspace $\operatorname{Pol}_{k,0}$. Therefore, $r_{k}=r_{k}-1-(-1)$ is orthogonal to $\operatorname{Pol}_{k,0}$ with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$. For $k=d$, the claim follows from 9.5(a). ∎
Lemma 9.3 implies the orthogonality of the residual polynomials. The approach of analysing the CG iterate by using the properties of zeros of orthogonal polynomials dates back to the seminal work of Nemirovskii [20] for deterministic inverse problems, which was further developed by Hanke [10] and Engl, Hanke and Neubauer [8].
Corollary 9.4.
The sequence $(r_{k})_{k=0,\dots,d}$ of residual polynomials is conjugate with respect to $AA^{\top}$ and $\langle{\cdot},{\cdot}\rangle_{Y}$, i.e.
[…]
where $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}>0$ for all $l=0,\dots,d-1$.
Proof.
Without loss of generality, assume $l<k$. Since $x\mapsto xr_{l}(x)\in\operatorname{Pol}_{l+1,0}\subset\operatorname{Pol}_{k,0}$, Lemma 9.3 yields $\langle{\mathsf{r_{k}}},{\mathsf{xr_{l}}}\rangle_{Y}=0$. For $l<d$, there is $i^{*}\in\{1,\dots,d\}$ such that $r_{l}(\tilde{\lambda}_{i^{*}}^{2})\neq 0$, since a polynomial of degree $l$ can have at most $l$ zeros. Hence, by (Y), $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}^{2}\geqslant\tilde{\lambda}_{i^{*}}^{2}r_{l}(\tilde{\lambda}_{i^{*}}^{2})^{2}\sum_{j:\,\lambda_{j}=\tilde{\lambda}_{i^{*}}}Y_{j}^{2}>0$, which concludes the proof. ∎
Note that $\lVert\mathsf{x^{1/2}r_{d}}\rVert_{Y}=0$. The following properties of the noninterpolated residual polynomials are well-known [8, Section 7.2 and Appendix A.2].
Lemma 9.5 (Properties of the residual polynomial $r_{k}$).
  1. (a)
    $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$, $k=1,\dots,d-1$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<\tilde{\lambda}_{d}^{2}<\dots<\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
  2. (b)
    $r_{k}$, $k=1,\dots,d$, can be written as
    $r_{k}(x)=\prod_{i=1}^{k}\Big{(}1-\frac{x}{x_{i,k}}\Big{)},\ x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\quad\text{with}\quad\lvert r_{k}^{\prime}(0)\rvert=\sum_{i=1}^{k}x_{i,k}^{-1}.$
  3. (c)
    $r_{k}$ is nonnegative, decreasing and convex on $[0,x_{1,k}]$ and log-concave on $[0,x_{1,k})$ for $k=1,\dots,d$.
  4. (d)
    For $k=1,\dots,d-1$, we have
    $x_{1,k+1}<x_{1,k}<x_{2,k+1}<\dots<x_{k,k+1}<x_{k,k}<x_{k+1,k+1}.$
  5. (e)
    $r_{k}$ is lower bounded by $r_{k+1}$ on $[0,x_{1,k+1}]$. In particular, $r_{k}(x)>r_{k+1}(x)$ for $0<x\leqslant x_{1,k+1}$ and $r_{k}(0)=r_{k+1}(0)=1$, $k=1,\dots,d-1$.
  6. (f)
    For $k=1,\dots,d$, $i=1,\dots,k$, we have
    $x_{k+1-i,k}\leqslant\lambda_{i}^{2}.$
Proof.
[…]
  1. 1.
    Corollary 9.4 shows that $(r_{k})_{k=0,\dots,d}$ is orthogonal in $L^{2}([0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\hat{\mu})$ with $\hat{\mu}\coloneqq\sum_{i=1}^{D}\lambda_{i}^{2}Y_{i}^{2}\delta_{\lambda_{i}^{2}}$ and $\lVert r_{k}\rVert_{L^{2}(\hat{\mu})}>0$ for $k=0,\dots,d-1$. The first part for $k<d$ is a well-known fact for orthogonal polynomials, see Szegő [24, Theorem 3.3.1]. Since $A$ has $d$ distinct nonzero singular values, we have $\lVert\mathsf{r_{d}}\rVert_{Y}^{2}=\min_{p\in\operatorname{Pol}_{d,1}}\lVert\mathsf{p}\rVert_{Y}^{2}=0$. By Assumption (Y), this implies $r_{d}(\tilde{\lambda}_{i}^{2})=0$, $i=1,\dots,d$, and we conclude the second part.
  2. 2.
[…]
    $r_{k}^{\prime}(x)=\sum_{i=1}^{k}\bigg{(}\Big{(}-\frac{1}{x_{i,k}}\Big{)}\prod_{\begin{subarray}{c}1\leqslant j\leqslant k\\ j\neq i\end{subarray}}\Big{(}1-\frac{x}{x_{j,k}}\Big{)}\bigg{)}.$
    This gives the claim by plugging in $x=0$.
  3. 3.
[…]
    $\log(r_{k}(x))=\sum_{i=1}^{k}\log\Big{(}1-\frac{x}{x_{i,k}}\Big{)}$
    is concave on $[0,x_{1,k})$ as a sum of concave functions.
  4. 4.
    Recall the definition of the Wronskian given in (9.1) and that $(r_{k})_{k=0,\dots,d}$ is orthogonal in $L^{2}([0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\hat{\mu})$ with $\lVert r_{k}\rVert_{L^{2}(\hat{\mu})}>0$ for $k=0,\dots,d-1$. For $k=1,\dots,d-1$, Lemma 9.2 yields
    $W(x;r_{k},r_{k+1})\neq 0,\quad x\in(0,\lVert AA^{\top}\rVert_{\mathrm{spec}}),$ (9.2)
[…]
    $x_{1,d}<x_{1,d-1}<x_{2,d}<x_{2,d-1}<\dots<x_{d-1,d}<x_{d-1,d-1}<x_{d,d},$
    and
    $x_{1,d-1}<x_{1,d}<x_{2,d-1}<x_{2,d}<\dots<x_{d-1,d-1}<x_{d-1,d}<x_{d,d}.$
    Assume the second case, where, in particular, $x_{d-1,d-1}\in(x_{d-2,d},x_{d-1,d})$. By considering the two cases whether $d$ is even or odd separately, one can show that $W(x_{d-1,d-1};r_{d-1},r_{d})>0$ and $W(x_{d,d};r_{d-1},r_{d})<0$. Since the Wronskian is continuous, it has a zero in $(x_{d-1,d-1},x_{d,d})\subset(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$. This contradicts (9.2), which shows the claim.
  5. 5.
[…]
    $r_{k+1}(x)=\prod_{i=1}^{k+1}\Big{(}1-\frac{x}{x_{i,k+1}}\Big{)}<\prod_{i=1}^{k}\Big{(}1-\frac{x}{x_{i,k}}\Big{)}=r_{k}(x)$
    for $0<x\leqslant x_{1,k+1}$, which shows the claim.
  6. 6.
[…]
    $x_{d+1-i,d}=\tilde{\lambda}_{i}^{2}\leqslant\lambda_{i}^{2},\quad i=1,\dots,d.$
[…]
    $x_{k,k}<x_{d,d}=\tilde{\lambda}_{1}^{2}=\lambda_{1}^{2}.$
    In particular, the claim holds for $k=1$. For $k=2,\dots,d-1$, fix $j\in\{1,\dots,d-1\}$ and assume that there exists an $l\in\{1,\dots,k-1\}$ such that $x_{l,k},x_{l+1,k}\in(\tilde{\lambda}_{j+1}^{2},\tilde{\lambda}_{j}^{2}]$. Let
    $f(x)\coloneqq\Big{(}\prod_{\begin{subarray}{c}i=1,\dots,k,\\ i\neq l,l+1\end{subarray}}(x-x_{i,k})^{2}\Big{)}(x-x_{l,k})(x-x_{l+1,k}),\quad x\in{\mathbb{R}}.$
    For $x\in\{\tilde{\lambda}_{1}^{2},\dots,\tilde{\lambda}_{d}^{2}\}$, there are two cases: if $x\leqslant\tilde{\lambda}_{j+1}^{2}$, then $x-x_{l+1,k}<x-x_{l,k}<0$; otherwise, if $x\geqslant\tilde{\lambda}_{j}^{2}$, then $0\leqslant x-x_{l+1,k}<x-x_{l,k}$. In both cases, we have
    $(x-x_{l,k})(x-x_{l+1,k})\geqslant 0.$
    Thus, $f(x)\geqslant 0$ for all $x\in\{\tilde{\lambda}_{1}^{2},\dots,\tilde{\lambda}_{d}^{2}\}$. In particular, $x_{i,k}<x_{k,k}<\tilde{\lambda}_{1}^{2}$, $i=1,\dots,k-1$, implies $f(\tilde{\lambda}_{1}^{2})>0$. Hence, under (Y),
    $\sum_{i=1}^{D}\lambda_{i}^{2}f(\lambda_{i}^{2})Y_{i}^{2}\geqslant\tilde{\lambda}_{1}^{2}f(\tilde{\lambda}_{1}^{2})\sum_{j:\,\lambda_{j}=\tilde{\lambda}_{1}}Y_{j}^{2}>0.$
    On the other hand, we can write
    $f(x)=c^{-1}r_{k}(x)\cdot s_{k-2}(x)\quad\text{with}\quad s_{k-2}(x)\coloneqq\prod_{\begin{subarray}{c}i=1,\dots,k,\\ i\neq l,l+1\end{subarray}}(x-x_{i,k})\in\operatorname{Pol}_{k-2},$
    where $c\coloneqq\prod_{i=1}^{k}-x_{i,k}^{-1}$. We deduce
    $\sum_{i=1}^{D}\lambda_{i}^{2}f(\lambda_{i}^{2})Y_{i}^{2}=\sum_{i=1}^{D}r_{k}(\lambda_{i}^{2})q_{k-1}(\lambda_{i}^{2})Y_{i}^{2}=\langle{\mathsf{r_{k}}},{\mathsf{q_{k-1}}}\rangle_{Y},$
    where $q_{k-1}(x)\coloneqq c^{-1}xs_{k-2}(x)\in\operatorname{Pol}_{k-1,0}$. By Lemma 9.3, we have $\langle{\mathsf{r_{k}}},{\mathsf{q_{k-1}}}\rangle_{Y}=0$, which yields a contradiction. We conclude that every interval $(\tilde{\lambda}_{j+1}^{2},\tilde{\lambda}_{j}^{2}]$, $j=1,\dots,d-1$, contains at most one zero of $r_{k}$. Since $x_{k,k}<\tilde{\lambda}_{1}^{2}$, this implies $x_{k+1-i,k}\leqslant\tilde{\lambda}_{i}^{2}\leqslant\lambda_{i}^{2}$ for $k=1,\dots,d$, $i=1,\dots,k$. ∎
The following result is originally due to Nemirovskii [20, (3.14)], who introduced the function $\varphi_{k}^{2}$, named $r_{k}$ therein. The proof of the first inequality can also be found in Hanke [10, (3.8)] and Engl, Hanke and Neubauer [8, (7.7)]. For the sake of completeness, we state it in our notation below. The second inequality is new, but follows easily from 9.5(b) and the definition of $\varphi_{k}$.
Lemma 9.6.
The residual polynomial satisfies
$\lVert\mathsf{r_{k}}\rVert_{Y}^{2}<\lVert\mathsf{\varphi_{k}}\rVert_{Y}^{2}\leqslant\lVert\mathsf{r_{k,<}^{1/2}}\rVert_{Y}^{2}$
for $k=1,\dots,d$, where $\varphi_{k}^{2}(x)\coloneqq\frac{x_{1,k}}{x_{1,k}-x}r_{k,<}^{2}(x)$, $x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]$. For $k=0$, $\lVert\mathsf{r_{k}}\rVert_{Y}^{2}\leqslant\lVert\mathsf{r_{k,<}^{1/2}}\rVert_{Y}^{2}$ continues to hold.
Proof.
For $k=0$, the claim is clear. Let $k=1,\dots,d$. We have
$\frac{x}{x-x_{1,k}}r_{k}(x)=-\frac{x}{x_{1,k}}\prod_{i=2}^{k}\Big{(}1-\frac{x}{x_{i,k}}\Big{)}\in\operatorname{Pol}_{k,0}.$
Lemma 9.3 yields $r_{k}\perp_{Y}\frac{x}{x-x_{1,k}}r_{k}$ such that
$\Big{\langle}{\mathsf{r_{k}}},{\mathsf{\frac{x}{x-x_{1,k}}r_{k}}}\Big{\rangle}_{Y,>}=\Big{\langle}{\mathsf{r_{k}}},{\mathsf{\frac{x}{x_{1,k}-x}r_{k}}}\Big{\rangle}_{Y,<}.$
Using
$\Big{\langle}{\mathsf{r_{k}}},{\mathsf{\frac{x}{x-x_{1,k}}r_{k}}}\Big{\rangle}_{Y,>}>\lVert\mathsf{r_{k}}\rVert_{Y,>}^{2},$
this implies
$\lVert\mathsf{r_{k}}\rVert_{Y}^{2}$ $=\lVert\mathsf{r_{k}}\rVert_{Y,<}^{2}+\lVert\mathsf{r_{k}}\rVert_{Y,>}^{2}$
$<\langle{\mathsf{r_{k}}},{\mathsf{r_{k}}}\rangle_{Y,<}+\Big{\langle}{\mathsf{r_{k}}},{\mathsf{\frac{x}{x_{1,k}-x}r_{k}}}\Big{\rangle}_{Y,<}$
$=\lVert\mathsf{\varphi_{k}}\rVert_{Y}^{2}.$
Note that $r_{k,<}(x)\leqslant 1-\frac{x}{x_{1,k}}=\frac{x_{1,k}-x}{x_{1,k}}$. Thus,
$\lVert\mathsf{\varphi_{k}}\rVert_{Y}^{2}=\Big{\lVert}\mathsf{\Big{(}\frac{x_{1,k}}{x_{1,k}-x}\Big{)}^{1/2}r_{k,<}^{1/2}r_{k,<}^{1/2}}\Big{\rVert}_{Y}^{2}\leqslant\lVert\mathsf{r_{k,<}^{1/2}}\rVert_{Y}^{2}.\qed$
9.3 Proofs for Section 4
Proof of Lemma 4.3.
For $t,s=0,\dots,d$, Lemma 9.3 yields
$R_{t\vee s}^{2}-\langle{\mathsf{r_{t}}},{\mathsf{r_{s}}}\rangle_{Y}=\langle{\mathsf{r_{t\vee s}-r_{t\wedge s}}},{\mathsf{r_{t\vee s}}}\rangle_{Y}=0,$
since $r_{t\vee s}-r_{t\wedge s}\in\operatorname{Pol}_{t\vee s,0}$. We conclude
$\lVert A(\widehat{f}_{t}-\widehat{f}_{s})\rVert^{2}$ $=\lVert\mathsf{r_{t}-r_{s}}\rVert_{Y}^{2}$
$=R_{t}^{2}+R_{s}^{2}-2\langle{\mathsf{r_{t}}},{\mathsf{r_{s}}}\rangle_{Y}$
$=R_{t}^{2}+R_{s}^{2}-2R_{t\vee s}^{2}$
$=\lvert R_{t}^{2}-R_{s}^{2}\rvert.$ (9.3)
This gives the claim for integer indices.
For $t=k+\alpha$ with $k=0,\dots,d-1$, $\alpha\in[0,1]$ and $0\leqslant s\leqslant k$, we have $r_{s}\in\operatorname{Pol}_{k,1}$, $r_{t}-r_{s}\in\operatorname{Pol}_{k+1,0}$ and $r_{k}-r_{s}\in\operatorname{Pol}_{k,0}$. Lemma 9.3 implies $\langle{\mathsf{r_{k+1}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}=0$, $\langle{\mathsf{r_{k}}},{\mathsf{r_{k}-r_{s}}}\rangle_{Y}=0$ and $\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}=R_{k+1}^{2}$. Thus, we obtain
$\langle{\mathsf{r_{t}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}$ $=\langle{(1-\alpha)\mathsf{r_{k}}+\alpha\mathsf{r_{k+1}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}$
$=(1-\alpha)\langle{\mathsf{r_{k}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}$
$=(1-\alpha)\langle{\mathsf{r_{k}}},{\mathsf{r_{k}-r_{s}}}\rangle_{Y}+(1-\alpha)\alpha\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}-r_{k}}}\rangle_{Y}$
$=(1-\alpha)\alpha(R_{k+1}^{2}-R_{k}^{2})$
$\leqslant 0.$
Similarly, for $t=k+\alpha$, $s=k+\beta$ with $0\leqslant\beta\leqslant\alpha\leqslant 1$, we deduce
$\langle{\mathsf{r_{t}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}$ $=(1-\alpha)\langle{\mathsf{r_{k}}},{\mathsf{r_{t}-r_{s}}}\rangle_{Y}$
$=(1-\alpha)\langle{\mathsf{r_{k}}},{(1-\alpha)\mathsf{r_{k}}+\alpha\mathsf{r_{k+1}}-(1-\beta)\mathsf{r_{k}}-\beta\mathsf{r_{k+1}}}\rangle_{Y}$
$=(1-\alpha)(\beta-\alpha)R_{k}^{2}+(1-\alpha)(\alpha-\beta)\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}$
$=(1-\alpha)(\alpha-\beta)\big{(}R_{k+1}^{2}-R_{k}^{2}\big{)}$
$\leqslant 0.$
We infer $\langle{\mathsf{r_{t}}},{\mathsf{r_{s}}}\rangle_{Y}\geqslant R_{t\vee s}^{2}$ for any $t,s\in[0,d]$. The second inequality follows similarly to (9.3) by applying $\langle{\mathsf{r_{t}}},{\mathsf{r_{s}}}\rangle_{Y}\geqslant R_{t\vee s}^{2}$. ∎
Proof of Lemma 4.5.
  1. 1.
    See 9.5(a).
  2. 2.
    Let $k=0$. 9.5(b) yields
    $r_{t}(x)=1-\alpha+\alpha\Big{(}1-\frac{x}{x_{1,1}}\Big{)}=1-\alpha\frac{x}{x_{1,1}},\quad x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],$
[…]
    $0<x_{1,k+1}<x_{1,k}<x_{2,k+1}<\dots<x_{k,k+1}<x_{k,k}<x_{k+1,k+1}\leqslant\lVert AA^{\top}\rVert_{\mathrm{spec}}.$
    For $x\in(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$, (9.2) implies
    $W(x;r_{t},r_{k+1})$ $=r_{t}(x)r_{k+1}^{\prime}(x)-r_{t}^{\prime}(x)r_{k+1}(x)$
    $=(1-\alpha)\big{(}r_{k}(x)r_{k+1}^{\prime}(x)-r_{k}^{\prime}(x)r_{k+1}(x)\big{)}$
    $=(1-\alpha)W(x;r_{k},r_{k+1})$
    $\neq 0,$
    and analogously $W(x;r_{t},r_{k})\neq 0$. Thus, by Lemma 9.1, any two consecutive zeros of $r_{k}$ or $r_{k+1}$ in $(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$ are separated by a zero of $r_{t}$ and vice versa, and the zeros of $r_{t}$ in $(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$ are simple. Now, $r_{t}(x)>0$ on $[0,x_{1,k+1}]$, since $r_{k}(x)>0$ and $r_{k+1}(x)\geqslant 0$ on $[0,x_{1,k+1}]$, and $r_{t}(x_{1,k})<0$, since $r_{k+1}(x)<0$ on $(x_{1,k+1},x_{1,k}]$. Hence, $x_{1,k+1}<x_{1,t}<x_{1,k}$. Thus, by the interlacing of $r_{t}$ and $r_{k}$ in $(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$, $r_{t}$ has at least $k$ simple zeros in $(0,\lVert AA^{\top}\rVert_{\mathrm{spec}})$. The $k+1$-th zero is also real and simple. Otherwise, by the complex conjugate root theorem, its complex conjugate would also be a zero, but this is not possible, since $r_{t}$ has degree $k+1$. We obtain by the interlacing
    $x_{1,t}<x_{1,k}<x_{2,t}<\dots<x_{k,t}<x_{k,k}<x_{k+1,t}$
    and
    $x_{1,k+1}<x_{1,t}<x_{2,k+1}<\dots<x_{k,k+1}<x_{k,t}.$
    By considering the two cases whether $k$ is even or odd separately, one can show that $r_{t}(x)\neq 0$ on $(x_{k,k},x_{k+1,k+1}]$ such that $x_{k+1,t}>x_{k+1,k+1}$. Combining this with the two chains of inequalities yields the claim.
  3. 3.
    Follows analogously to 9.5(b).
  4. 4.
    Follows analogously to 9.5(c).
  5. 5.
    The definitions of $r_{t_{1}}$ and $r_{t_{2}}$ and (9.2) yield
    $W(x;r_{t_{1}},r_{t_{2}})=(\alpha_{2}-\alpha_{1})W(x;r_{k},r_{k+1})\neq 0,\quad x\in(0,\lVert AA^{\top}\rVert_{\mathrm{spec}}).$
[…]
    $r_{t_{1}}(x)=r_{t_{2}}(x)+(\alpha_{2}-\alpha_{1})(r_{k}(x)-r_{k+1}(x))>r_{t_{2}}(x),$
    since $r_{k+1}(x)<r_{k}(x)$ for all $x\in(0,x_{1,k+1}]$ by 9.5(e), $r_{k}(x)\geqslant 0$ for $x\in[0,x_{1,k}]$ and $r_{k+1}(x)<0$ for $x\in(x_{1,k+1},x_{1,k}]$. This implies $x_{1,t_{2}}<x_{1,t_{1}}$. Let $x>x_{k+1,k+1}$. For $k$ even we obtain $r_{k}(x)>0$ and $r_{k+1}(x)<0$ such that $r_{t_{1}}(x)>r_{t_{2}}(x)$. Since $r_{t_{j}}(x_{k+1,k+1})>0$, $j=1,2$, this yields $x_{k+1,t_{2}}<x_{k+1,t_{1}}$. In the same way, for $k$ odd we deduce $r_{t_{1}}(x)<r_{t_{2}}(x)$ and $r_{t_{j}}(x_{k+1,k+1})<0$, $j=1,2$, implying $x_{k+1,t_{2}}<x_{k+1,t_{1}}$.
  6. 6.
[…]
  7. 7.
    See 9.5(f). ∎
Proof of Proposition 4.6.
For integer $t=0,\dots,d$, this was proved in Lemma 9.6. Now, consider the case $t=k+\alpha$ with $k=0,\dots,d-1$, $\alpha\in(0,1)$. By Lemma 9.3, we have $r_{k}\perp_{Y}\operatorname{Pol}_{k,0}$ and $r_{k+1}\perp_{Y}\operatorname{Pol}_{k+1,0}\supset\operatorname{Pol}_{k,0}$ such that $r_{t}\perp_{Y}\operatorname{Pol}_{k,0}$. Using in addition that $1-r_{k}\in\operatorname{Pol}_{k,0}$, $1-r_{k+1}\in\operatorname{Pol}_{k+1,0}$ and $r_{k+1}-r_{k}\in\operatorname{Pol}_{k+1,0}$, we obtain
$\langle{\mathsf{r_{t}}},{\mathsf{1}}\rangle_{Y}=(1-\alpha)\langle{\mathsf{r_{k}}},{\mathsf{1}}\rangle_{Y}+\alpha\langle{\mathsf{r_{k+1}}},{\mathsf{1}}\rangle_{Y}=(1-\alpha)R_{k}^{2}+\alpha R_{k+1}^{2}$
and
$\langle{\mathsf{r_{t}}},{\mathsf{r_{k+1}-1}}\rangle_{Y}$ $=(1-\alpha)\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}-1}}\rangle_{Y}$
$=(1-\alpha)\langle{\mathsf{r_{k}-r_{k+1}}},{\mathsf{r_{k+1}-1}}\rangle_{Y}$
$=-(1-\alpha)\langle{\mathsf{r_{k}-r_{k+1}}},{\mathsf{1}}\rangle_{Y}$
$=-(1-\alpha)(R_{k}^{2}-R_{k+1}^{2}).$
We deduce the orthogonality relation
$\langle{\mathsf{r_{t}}},{\mathsf{r_{k+1}-1+\zeta}}\rangle_{Y}=-(1-\alpha)(R_{k}^{2}-R_{k+1}^{2})+\zeta\big{(}(1-\alpha)R_{k}^{2}+\alpha R_{k+1}^{2}\big{)}=0$ (9.4)
with
$\zeta\coloneqq\tfrac{(1-\alpha)(R_{k}^{2}-R_{k+1}^{2})}{(1-\alpha)R_{k}^{2}+\alpha R_{k+1}^{2}}\leqslant 1.$
Put
$\tilde{x}_{1,t}\coloneqq\zeta x_{1,{k+1}}\prod_{i=2}^{k+1}\tfrac{x_{i,k+1}}{x_{i,t}},\quad\tilde{r}_{t}\coloneqq\tfrac{\tilde{x}_{1,t}-x}{x_{1,t}-x}r_{t},$
and note the order $0\leqslant\tilde{x}_{1,t}<x_{1,k+1}<x_{1,t}$ due to $x_{i,t}>x_{i,k+1}$ by 4.5(b). With some $q_{k},\tilde{q}_{k},\check{q}_{k}\in\operatorname{Pol}_{k,0}$, $\tilde{p}_{k-1}\in\operatorname{Pol}_{k-1,1}$, we have
$\tilde{r}_{t}(x)$ $=\frac{\tilde{x}_{1,t}-x}{x_{1,t}-x}\prod_{i=1}^{k+1}\Big{(}1-\frac{x}{x_{i,t}}\Big{)}$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}\prod_{i=2}^{k+1}\Big{(}1-\frac{x}{x_{i,t}}\Big{)}-\frac{x}{x_{1,t}}\prod_{i=2}^{k+1}\Big{(}1-\frac{x}{x_{i,t}}\Big{)}$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}+\check{q}_{k}(x)-\frac{x}{x_{1,t}}\bigg{(}x^{k}\prod_{i=2}^{k+1}\Big{(}-\frac{1}{x_{i,t}}\Big{)}+\tilde{p}_{k-1}(x)\bigg{)}$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}+\tilde{q}_{k}(x)-\frac{x^{k+1}}{x_{1,t}}\prod_{i=2}^{k+1}\Big{(}-\frac{1}{x_{i,t}}\Big{)}$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}+\tilde{q}_{k}(x)+\frac{x^{k+1}}{x_{1,t}}\tilde{x}_{1,t}\zeta^{-1}\prod_{i=1}^{k+1}\Big{(}-\frac{1}{x_{i,k+1}}\Big{)}$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}\zeta^{-1}\bigg{(}\zeta+\prod_{i=1}^{k+1}\Big{(}-\frac{x}{x_{i,k+1}}\Big{)}\bigg{)}+\tilde{q}_{k}(x)$
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}\zeta^{-1}\big{(}\zeta+r_{k+1}(x)-1-q_{k}(x)\big{)}+\tilde{q}_{k}(x).$
The orthogonality relation (9.4) together with $r_{t}\perp_{Y}\operatorname{Pol}_{k,0}$ therefore gives $\tilde{r}_{t}\perp_{Y}r_{t}$. Consequently, $\langle{\mathsf{\tilde{r}_{t}}},{\mathsf{r_{t}}}\rangle_{Y,>}=-\langle{\mathsf{\tilde{r}_{t}}},{\mathsf{r_{t}}}\rangle_{Y,<}$ holds and thus
$\lVert\mathsf{r_{t}}\rVert_{Y}^{2}$ $=\lVert\mathsf{r_{t,<}}\rVert_{Y}^{2}+\lVert\mathsf{r_{t,>}}\rVert_{Y}^{2}\leqslant\lVert\mathsf{r_{t,<}}\rVert_{Y}^{2}+\langle{\mathsf{r_{t}}},{\mathsf{(\tfrac{x-\tilde{x}_{1,t}}{x-x_{1,t}})r_{t}}}\rangle_{Y,>}=\lVert\mathsf{r_{t,<}}\rVert_{Y}^{2}+\langle{\mathsf{r_{t}}},{\mathsf{\tilde{r}_{t}}}\rangle_{Y,>}$
$=\lVert\mathsf{r_{t,<}}\rVert_{Y}^{2}-\langle{\mathsf{r_{t}}},{\mathsf{\tilde{r}_{t}}}\rangle_{Y,<}=\langle{\mathsf{r_{t}(r_{t}-\tilde{r}_{t})}},{\mathsf{1}}\rangle_{Y,<}=\lVert\mathsf{(r_{t,<}^{2}-r_{t,<}\tilde{r}_{t,<})^{1/2}}\rVert_{Y}^{2}.$
From $r_{t,<}(x)\leqslant(1-x_{1,t}^{-1}x)_{+}$ by 4.5(c), we infer
$(r_{t,<}^{2}-r_{t,<}\tilde{r}_{t,<})(x)$ $=r_{t,<}^{2}(x)\Big{(}1-\frac{\tilde{x}_{1,t}-x}{x_{1,t}-x}\Big{)}=r_{t,<}^{2}(x)\frac{x_{1,t}-\tilde{x}_{1,t}}{x_{1,t}-x}$
$\leqslant r_{t,<}(x)\frac{x_{1,t}-\tilde{x}_{1,t}}{x_{1,t}}\leqslant r_{t,<}(x).$
This gives the result. ∎
Proof of Lemma 4.7.
For $t=0$, the claim is clear. Let $t\in(0,d]$. Recall that $r_{t}^{\prime}(0)<0$ and that $r_{t}$ is convex and nonnegative on $[0,x_{1,t}]$ by 4.5(d). This implies
$(1-\lvert r_{t}^{\prime}(0)\rvert x)_{+}\leqslant r_{t}(x)=\prod_{i=1}^{\lceil t\rceil}\Big{(}1-\frac{x}{x_{i,t}}\Big{)}\leqslant 1-\frac{x}{x_{1,t}},\quad x\in[0,x_{1,t}].$
By the log-concavity in 4.5(d), note for $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in(0,1]$,
$\log(r_{t}(x))\leqslant\log(r_{t}(0))+x\Big{(}\frac{d}{dy}\log(r_{t}(y))\Big{)}\Big{|}_{y=0}=r_{t}^{\prime}(0)x,\quad x\in[0,x_{1,t}).$
We conclude
$(1-\lvert r_{t}^{\prime}(0)\rvert x)_{+}\leqslant r_{t}(x)\leqslant\exp(r_{t}^{\prime}(0)x)=\exp(-\lvert r_{t}^{\prime}(0)\rvert x),\quad x\in[0,x_{1,t}].\qed$
Proof of Corollary 4.8.
Note that $\lvert r_{t}^{\prime}(0)\rvert=\sum_{i=1}^{\lceil t\rceil}x_{i,t}^{-1}>0$ by 4.5(c). We have
$\sup_{x\geqslant 0}e^{-\lvert r_{t}^{\prime}(0)\rvert x}x^{2\mu+1}=\Big{(}\frac{2\mu+1}{e\lvert r_{t}^{\prime}(0)\rvert}\Big{)}^{2\mu+1}.$
Under the source condition $\lVert g\rVert_{\mu+1/2}\leqslant R$, we deduce by Lemma 4.7
$\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}$ $\leqslant\lVert\mathsf{\exp(-\lvert r_{t}^{\prime}(0)\rvert x/2)}g\rVert^{2}$
$=\sum_{i=1}^{D}\big{(}\exp(-\lvert r_{t}^{\prime}(0)\rvert\lambda_{i}^{2})\lambda_{i}^{4\mu+2}\big{)}\lambda_{i}^{-(4\mu+2)}g_{i}^{2}$
$\leqslant\Big{(}\frac{2\mu+1}{e\lvert r_{t}^{\prime}(0)\rvert}\Big{)}^{2\mu+1}\lVert g\rVert_{\mu}^{2}$
$\leqslant R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{t}^{\prime}(0)\rvert^{-2\mu-1}.\qed$
9.4 Proofs for Section 5
Proof of Lemma 5.5.
Noting that $r_{t,<}\leqslant 1$, we have
$\lVert\mathsf{r_{t,<}}Y\rVert^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}=-\langle{\mathsf{r_{t,<}(1-r_{t,<})}Y},{Y}\rangle=-\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}Y\rVert^{2},$
and analogously
$\lVert\mathsf{r_{t,<}}g\rVert^{2}-\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}=-\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}g\rVert^{2}.$
Hence, the approximation error can be expressed as
$A_{t,\lambda}=$ $\lVert\mathsf{r_{t,<}}g\rVert^{2}+\lVert\mathsf{r_{t,>}}Y\rVert^{2}$
$-\big{(}\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}Y\rVert^{2}-\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}g\rVert^{2}\big{)}$
with
$\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}Y\rVert^{2}$
$=\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}g\rVert^{2}+\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}\xi\rVert^{2}+2\langle{\mathsf{r_{t,<}}g},{(\mathsf{1-r_{t,<}})\xi}\rangle.$
This implies
$\lVert\mathsf{r_{t,<}}g\rVert^{2}\leqslant A_{t,\lambda}+\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}\xi\rVert^{2}+2\langle{\mathsf{r_{t,<}}g},{(\mathsf{1-r_{t,<}})\xi}\rangle.$
By the Cauchy–Schwarz inequality and the definition of $S_{t,\lambda}$, we deduce
$\lVert\mathsf{r_{t,<}}g\rVert^{2}$ $\leqslant A_{t,\lambda}+\lVert\mathsf{(r_{t,<}(1-r_{t,<}))^{1/2}}\xi\rVert^{2}+2\lVert\mathsf{r_{t,<}}g\rVert\lVert(\mathsf{1-r_{t,<}})\xi\rVert$
$\leqslant A_{t,\lambda}+S_{t,\lambda}+2S_{t,\lambda}^{1/2}\lVert\mathsf{r_{t,<}}g\rVert.$
For $x,A,B\geqslant 0$, $x^{2}\leqslant Ax+B$ implies $x^{2}\leqslant A^{2}+2B$, which is immediate for $x\leqslant A$ and follows for $x>A$ from $(x+A)(x-A)\leqslant 2x(x-A)\leqslant 2B$. Noting that $B\coloneqq A_{t,\lambda}+S_{t,\lambda}\geqslant\tfrac{1}{2}\lVert\widehat{g}_{t}-g\rVert^{2}\geqslant 0$ by Proposition 5.2, we obtain $\lVert\mathsf{r_{t,<}}g\rVert^{2}\leqslant 6S_{t,\lambda}+2A_{t,\lambda}$. ∎
Proof of Proposition 5.11.
We prove the claim analogously to Johnstone [15, Proposition 4.23]. Without loss of generality, assume that $v_{1},\dots,v_{D}$ are the first $D$ canonical unit vectors. Let $\Theta^{m}(\eta)$, $m=1,\dots,D$, for $\eta\geqslant 0$ denote the set of $f\in{\mathbb{R}}^{D}$ that satisfy $\lvert f_{i}\rvert\leqslant\eta$ for $i=1,\dots,m$ and $f_{i}=0$ for $i=m+1,\dots,D$. Define the loss
$L_{A}(\widehat{f},f)\coloneqq\lVert A(f-\widehat{f})\rVert^{2}=\sum_{i=1}^{D}\lambda_{i}^{2}(f_{i}-\widehat{f}_{i})^{2}=\sum_{i=1}^{D}L_{A,i}(\widehat{f}_{i},f_{i}),$
where $L_{A,i}(\widehat{f}_{i},f_{i})\coloneqq\lambda_{i}^{2}(f_{i}-\widehat{f}_{i})^{2}$, $i=1,\dots,D$. The function $\widehat{f}_{i}\mapsto L_{A,i}(\widehat{f}_{i},f_{i})$ is convex and lower semicontinuous for any $f_{i}$. Thus, we can apply Johnstone [15, Proposition 4.16] and obtain
$\inf_{\widehat{f}}\sup_{f\in\Theta^{m}(\eta)}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A(f-\widehat{f})\rVert^{2}\big{]}$ $=\sum_{i=1}^{m}\Big{(}\inf_{\widehat{f}_{i}}\sup_{f_{i}\in[-\eta,\eta]}\operatorname{{\mathbb{E}}}_{f}\big{[}\lambda_{i}^{2}(f_{i}-\widehat{f}_{i})^{2}\big{]}\Big{)}$
$=\sum_{i=1}^{m}\lambda_{i}^{2}\Big{(}\inf_{\widehat{f}_{i}}\sup_{f_{i}\in[-\eta,\eta]}\operatorname{{\mathbb{E}}}_{f}\big{[}(f_{i}-\widehat{f}_{i})^{2}\big{]}\Big{)}$
$=\gamma\sum_{i=1}^{m}\lambda_{i}^{2}(\eta^{2}\wedge\delta^{2}\lambda_{i}^{-2})$
with $\gamma\in[2/5,1]$, where we used Johnstone [15, Eq. (4.40)] for the last equality. Setting $\eta\coloneqq\delta m^{p}$, under (PSD($p$,$c_{A}$$\infty$)), we deduce
$\inf_{\widehat{f}}\sup_{f\in\Theta^{m}(\delta m^{p})}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A(f-\widehat{f})\rVert^{2}\big{]}=\gamma\sum_{i=1}^{m}\lambda_{i}^{2}\eta^{2}\wedge\delta^{2}\geqslant(c_{A}\wedge 1)^{2}\gamma m\delta^{2}.$
We have
$\sum_{i=1}^{m}\lambda_{i}^{-4\mu}(\delta m^{p})^{2}\leqslant\delta^{2}m^{2p}c_{A}^{-4\mu}\sum_{i=1}^{m}i^{4\mu p}\leqslant c_{A}^{-4\mu}\delta^{2}m^{4\mu p+2p+1}.$
Hence, for $\Theta^{m}(\delta m^{p})$ to be contained in the class of signals $f$ satisfying $\lVert f\rVert_{\mu}\leqslant R$, it suffices that $c_{A}^{-4\mu}\delta^{2}m^{4\mu p+2p+1}\leqslant R^{2}$. Let $m_{1}$ be the largest integer such that $c_{A}^{-4\mu}m_{1}^{4\mu p+2p+1}\leqslant R^{2}/\delta^{2}$ and $m_{0}\in{\mathbb{R}}$ the solution to $c_{A}^{-4\mu}m_{0}^{4\mu p+2p+1}=R^{2}/\delta^{2}$. If $m_{0}\geqslant 4$, then $m_{1}/m_{0}\geqslant 1/2$, which can be shown by contradiction. For $\delta/R\leqslant d_{\mu,p,c_{A}}$, we conclude
$\inf_{\widehat{f}}\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A(f-\widehat{f})\rVert^{2}\big{]}$
$\geqslant\inf_{\widehat{f}}\sup_{f\in\Theta^{m}(\delta m_{1}^{p})}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A(f-\widehat{f})\rVert^{2}\big{]}$
$\geqslant(c_{A}\wedge 1)^{2}\gamma m_{1}\delta^{2}\geqslant(c_{A}\wedge 1)^{2}\gamma\frac{m_{0}}{2}\delta^{2}$
$=(c_{A}\wedge 1)^{2}c_{A}^{4\mu/(4\mu p+2p+1)}\frac{\gamma}{2}R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}.$
This gives the claim. ∎
9.5 Proofs for Section 7
Proof of Proposition 7.8.
  1. 1.
    If $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, then by Lemma 6.5
    $\lvert r_{\tau}^{\prime}(0)\rvert\lesssim_{p,c_{A}}(\delta^{-2}S_{\tau,\lambda})^{2p}+(\log d)^{2p}+z^{2p}$
    with probability at least $1-e^{-z}$ and we conclude by Corollary 7.3.
    If $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, then we obtain the rough bound
    $\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\leqslant\lambda_{D}^{-2}\lVert A(\widehat{f}_{\tau}-f^{\dagger})\rVert^{2}\leqslant 4\lambda_{D}^{-2}M_{\tau,\lambda},$ (9.5)
    using the prediction error bound from Proposition 5.2. Moreover, we then have by Lemma 4.7
    $S_{\tau,\lambda}\geqslant\lVert\mathsf{(1-{\exp(-\lambda_{D}^{-2}x)})^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\lVert\xi\rVert^{2}.$
    The lower $\chi^{2}$-concentration bound of Laurent and Massart [16, Lemma 1] gives with probability at least $1-e^{-z}$
    $\delta^{-2}\lVert\xi\rVert^{2}\geqslant D-2\sqrt{D}\sqrt{z},\quad\text{i.e.}\quad(\sqrt{D}-\sqrt{z})^{2}\leqslant\delta^{-2}\lVert\xi\rVert^{2}+z.$
    This shows by PSD($p$,$c_{A}$,$C_{A}$)
    $\lambda_{D}^{-2}\lesssim_{c_{A}}D^{2p}\leqslant(\delta^{-1}\lVert\xi\rVert+2\sqrt{z})^{4p}\lesssim_{p}(\delta^{-2}S_{\tau,\lambda})^{2p}+z^{2p},$
    which inserted into (9.5) yields an even sharper bound than claimed.
  2. 2.
    Fix $z>0$ and note $M_{\tau,\lambda}=A_{\tau,\lambda}\wedge S_{\tau,\lambda}+\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$. By 5.3(d), the source condition on $f^{\dagger}$ and (5.4) we have
    $A_{\tau,\lambda}\wedge S_{\tau,\lambda}$ $\leqslant\big{(}R^{2}(\mu+1/2)^{2\mu+1}\lvert r_{\tau}^{\prime}(0)\rvert^{-2\mu-1}\wedge\lVert(\mathsf{(\lvert r_{\tau}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}\big{)}$
    $\leqslant\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{)}$
    $\leqslant\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+\delta^{2}\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}\Big{)}.$ (9.6)
    Choosing $\rho\thicksim(R\delta^{-1})^{4p/(4\mu p+2p+1)}$ gives the right order of the first summand. For the second summand, we obtain similarly to (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)^{2}\leqslant C_{A}^{4}\sum_{i=1}^{D}\rho^{2}i^{-4p}\wedge 1\leqslant C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}.$
    The $\chi^{2}$-concentration bound of Laurent and Massart [16, Lemma 1] gives
    $\operatorname{{\mathbb{P}}}\Big{(}\textstyle\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)(Z_{i}^{2}-1)\geqslant 2C_{A}^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\Big{)}\leqslant e^{-z}.$
    From $2C_{A}^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\leqslant C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z$ we deduce with (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}\leqslant\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)+C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z\leqslant(1\vee C_{A}^{4})\tfrac{4p}{2p-1}\rho^{1/(2p)}+3z$
    with probability at least $1-e^{-z}$. Plugging into (9.6) for our choice of $\rho$ yields the claim.
  3. 3.
    Fix $z\geqslant 1$ and put $\Delta:=\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$ for short. Using (6.1), the identity $R_{\tau}^{2}=\kappa$ and the assumption on $\kappa$, we bound
    $\Delta$ $\leqslant\big{\lvert}\lVert\xi\rVert^{2}-\delta^{2}D\big{\rvert}+\lvert\kappa-\delta^{2}D\rvert+2\lvert\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert$
    $\leqslant\big{\lvert}\lVert\xi\rVert^{2}-\delta^{2}D\big{\rvert}+2\lvert\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert+C_{\kappa}\delta^{2}\sqrt{D}.$ (9.7)
    By the $\chi^{2}$-concentration bound of Laurent and Massart [16, Lemma 1], we have
    $\big{\lvert}\lVert\xi\rVert^{2}-\delta^{2}D\big{\rvert}\leqslant 2\delta^{2}\sqrt{D}\sqrt{z}+2\delta^{2}z$ (9.8)
    with probability at least $1-2e^{-z}$.
    Let us first assume $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$. Lemma 5.5 yields $\lVert\mathsf{r_{\tau,<}}g\rVert^{2}\leqslant 8M_{\tau,\lambda}$. We then obtain by (6.4), (9.7) and (9.8)
    $\delta^{-2}\Delta$
    $\lesssim_{p,c_{A},C_{A},C_{\kappa}}(\delta^{-2}M_{\tau,\lambda})^{1/2}\big{(}(\delta^{-2}S_{\tau,\lambda})^{p/(2p+1)}+\sqrt{\log d}+\sqrt{z}\big{)}+\sqrt{D}\sqrt{z}+z$
    $\lesssim_{p,c_{A},C_{A},C_{\kappa}}(\delta^{-2}M_{\tau,\lambda})^{(4p+1)/(4p+2)}+(\delta^{-2}M_{\tau,\lambda})^{1/2}\big{(}\sqrt{\log d}+\sqrt{z}\big{)}+\sqrt{D}\sqrt{z}+z$
    with probability at least $1-4e^{-z}$. Choosing as conjugate exponents $\tilde{p}\coloneqq(4p+1)/(2p+1)$ and $\tilde{q}\coloneqq(4p+1)/(2p)$, Young’s inequality implies
    $(\delta^{-2}M_{\tau,\lambda})^{1/2}(\sqrt{\log d}+\sqrt{z})$
    $\leqslant 2(\delta^{-2}M_{\tau,\lambda})^{(4p+1)/(4p+2)}+(\log d)^{(4p+1)/(4p)}+z^{(4p+1)/(4p)}.$
    We conclude
    $\delta^{-2}\Delta\lesssim_{p,c_{A},C_{A},C_{\kappa}}(\delta^{-2}M_{\tau,\lambda})^{(4p+1)/(4p+2)}+\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}$ (9.9)
[…]
    $\delta^{-2}M_{\tau,\lambda}\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}$ $(R\delta^{-1})^{2/(4\mu p+2p+1)}$
    $+(\delta^{-2}M_{\tau,\lambda})^{(4p+1)/(4p+2)}+\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}$
    with probability at least $1-5e^{-z}$. Due to $R\delta^{-1}\geqslant 1$ and $(4p+1)/(4p+2)<1$, we arrive for $z\geqslant 1$ at
    $\delta^{-2}M_{\tau,\lambda}\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}(R\delta^{-1})^{2/(4\mu p+2p+1)}+\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}$ (9.10)
    with probability at least $1-5e^{-z}$.
[…]
    $2\lvert\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert\leqslant 2\lVert\xi\rVert\lVert\mathsf{r_{\tau,<}}g\rVert\leqslant 2(1-e^{-1})^{-1/2}S_{\tau,\lambda}^{1/2}\lVert\mathsf{\exp(-\lambda_{D}^{-2}x)}g\rVert.$
    In view of
    $\sup_{x>0}\exp(-2\lambda_{D}^{-2}x)x^{2\mu+1}=\big{(}(\mu+1/2)e^{-1}\big{)}^{2\mu+1}\lambda_{D}^{4\mu+2}$
    and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we obtain
    $\lVert\mathsf{\exp(-\lambda_{D}^{-2}x)}g\rVert\leqslant\big{(}(\mu+1/2)e^{-1}\big{)}^{\mu+1/2}\lambda_{D}^{2\mu+1}R\lesssim_{\mu,C_{A}}RD^{-2\mu p-p}.$
    Combining with (9.7) and (9.8) yields with probability at least $1-2e^{-z}$
    $\delta^{-2}\Delta\lesssim_{\mu,C_{A},C_{\kappa}}(\delta^{-2}M_{\tau,\lambda})^{1/2}R\delta^{-1}D^{-2\mu p-p}+\sqrt{D}\sqrt{z}+z.$
[…]
    $\delta^{-2}M_{\tau,\lambda}\lesssim_{\mu,p,C_{A},C_{\kappa}}(R\delta^{-1})^{2/(4\mu p+2p+1)}+(\delta^{-2}M_{\tau,\lambda})^{1/2}R\delta^{-1}D^{-2\mu p-p}+\sqrt{D}\sqrt{z}+z$
    and thus for $z\geqslant 1$
    $\delta^{-2}M_{\tau,\lambda}\lesssim_{\mu,p,C_{A},C_{\kappa}}(R\delta^{-1})^{2/(4\mu p+2p+1)}+R^{2}\delta^{-2}D^{-4\mu p-2p}+\sqrt{D}\sqrt{z}+z.$
    From the assumption $D\gtrsim(R\delta^{-1})^{2/(4\mu p+2p+1)}$ we thus derive (9.10) again. ∎
Proof of Theorem 7.10.
Fix $z\geqslant 1$. By 7.8(a), the inequality
$\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}$ $\lesssim_{p,c_{A}}$ $R^{2/(2\mu+1)}M_{\tau,\lambda}^{2\mu/(2\mu+1)}+\delta^{-4p}S_{\tau,\lambda}^{2p}M_{\tau,\lambda}$
$+(\log d)^{2p}M_{\tau,\lambda}+z^{2p}M_{\tau,\lambda}\eqqcolon T_{1}+T_{2}+T_{3}+T_{4}$
holds true with probability at least $1-e^{-z}$. We control the summands using 7.8(c). The following bounds hold simultaneously with probability at least $1-5e^{-z}$:
We have
$T_{1}\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}$ $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$
$+R^{2/(2\mu+1)}\delta^{4\mu/(2\mu+1)}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2\mu/(2\mu+1)}.$
Young’s inequality with the conjugate exponents $\tilde{p}\coloneqq(2\mu+1)(2p+1)/(4\mu p+2p+1)$ and $\tilde{q}\coloneqq(2\mu+1)(2p+1)/(2\mu)$ yields
$\big{(}R^{2}\delta^{-2}\big{)}^{1/(2\mu+1)}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2\mu/(2\mu+1)}$
$\leqslant\big{(}R^{2}\delta^{-2}\big{)}^{(2p+1)/(4\mu p+2p+1)}+\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2p+1}$
such that
$T_{1}\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)+\delta^{2}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2p+1}.$
Furthermore,
$T_{2}$ $\lesssim\delta^{-4p}M_{\tau,\lambda}^{2p+1}$
$\lesssim_{\mu,p,c_{A},C_{A},C_{\kappa}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)+\delta^{2}\big{(}\sqrt{D}\sqrt{z}+z^{(4p+1)/(4p)}\big{)}^{2p+1}.$
By Young’s inequality with $\tilde{p}\coloneqq(2p+1)/(2p)$ and $\tilde{q}\coloneqq 2p+1$, we deduce
$T_{3}=\delta^{-4p}\big{(}\delta^{2}\log d\big{)}^{2p}M_{\tau,\lambda}\leqslant\delta^{-4p}M_{\tau,\lambda}^{2p+1}+\delta^{2}(\log d)^{2p+1}$
and
$T_{4}=\delta^{-4p}(\delta^{2}z)^{2p}M_{\tau,\lambda}\leqslant\delta^{-4p}M_{\tau,\lambda}^{2p+1}+\delta^{2}z^{2p+1}.$
We conclude the asserted high probability bound by combining all inequalities. By Lemma 9.7, the expectation bound follows directly from the high probability bound.
If $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, then $\delta^{2}D^{p+1/2}\lesssim\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ holds and we obtain minimax adaptivity for the given range of regularity parameters $\mu$ by Remark 7.7. ∎
Proof of Proposition 7.12.
Proposition 5.14 implies
$c\delta=R_{\tau}\leqslant\delta+R(\mu+1/2)^{\mu+1/2}\lvert r_{\tau}^{\prime}(0)\rvert^{-\mu-1/2}.$
Since $c>1$, this yields $\lvert r_{\tau}^{\prime}(0)\rvert\lesssim_{\mu,c}(R^{2}\delta^{-2})^{1/(2\mu+1)}$. By combining (7.1)–(7.3), we have
$\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}\lesssim\lVert\mathsf{r_{\tau,<}}g\rVert^{4\mu/(2\mu+1)}R^{2/(2\mu+1)}+\lvert r_{\tau}^{\prime}(0)\rvert S_{\tau,\lambda}+\lvert r_{\tau}^{\prime}(0)\rvert\lVert A(\widehat{f}_{\tau}-f)\rVert^{2}.$
Noting that $\lVert\mathsf{r_{\tau,<}}g\rVert\vee\lVert A(\widehat{f}_{\tau}-f)\rVert\leqslant R_{\tau}+\lVert\xi\rVert\leqslant(c+1)\delta$ and $S_{\tau,\lambda}\leqslant\delta^{2}$ under (DN), we obtain the claim by plugging in the bound on $\lvert r_{\tau}^{\prime}(0)\rvert$. ∎
9.6 Auxiliary result
Lemma 9.7.
Let $X$ be a real-valued random variable and $p\geqslant 1$ such that
$\operatorname{{\mathbb{P}}}(X\geqslant z^{p})\leqslant Ce^{-z}$
for all $z>c$ with absolute constants $c,C\geqslant 0$. Then
$\operatorname{{\mathbb{E}}}[X_{+}]\leqslant c^{p}+Cp\operatorname{{\mathbb{E}}}[Z^{p-1}],$
where $Z\sim\operatorname{Exp}(1)$ is an exponentially distributed random variable with parameter $1$.
Proof.
We have
[…]
Acknowledgements
Financial support by the Deutsche Forschungsgemeinschaft (DFG) through the research unit FOR 5381 Mathematical Statistics in the Information Age – Statistical Efficiency and Computational Tractability is gratefully acknowledged.